### v1

# Temporal reasoning for timeline summarisation in social media

Abstract
This paper explores whether enhancing temporal reasoning capabilities in Large Language Models (LLMs) can improve the quality of timeline summarization, the task of summarising long texts containing sequences of events, particularly social media threads . We introduce NarrativeReason, a novel dataset focused on temporal relationships among sequential events within narratives, distinguishing it from existing temporal reasoning datasets that primarily address pair-wise event relationships. Our approach then combines temporal reasoning with timeline summarization through a knowledge distillation framework, where we first fine-tune a teacher model on temporal reasoning tasks and then distill this knowledge into a student model while simultaneously training it for the task of timeline summarization. Experimental results demonstrate that our model achieves superior performance on mental health-related timeline summarization tasks, which involve long social media threads with repetitions of events and a mix of emotions, highlighting the importance of leveraging temporal reasoning to improve timeline summarisation.
1 Introduction
Timeline summarization organizes and presents a sequence of events in a coherent and concise manner Steen and Markert (2019); Li et al. (2021); Hu et al. (2024). It involves extracting event-related timelines and then summarising them Hu et al. (2024); Rajaby Faghihi et al. (2022). Researchers generally create event graphs Li et al. (2021) or cluster event related timelines Hu et al. (2024) to identify relevant events. Recent work Song et al. (2024) has introduced the challenging task of social media timeline summarisation, especially in the context of capturing fluctuations in individuals’ state of mind as reflected in posts shared online over time. In these posts, numerous events may occur without explicit timestamps, requiring contextual inference to determine their chronological sequence. Moreover, mental health-related events are not easy to identify: they can be connected to an individual’s emotions, interpersonal interactions, and the entire timeline is necessary to provide enough contextSong et al. (2024). It is particularly challenging to identify events pertaining to psychological states and to extract these from posts. When generating mental health related summaries from longitudinal posts, models need to understand related events and maintain temporal consistency to make inferences. This raises the question of whether temporal reasoning can be leveraged to enhance the quality of complex timeline summaries.
Temporal reasoning involves understanding and processing temporal information in text to deduce time-based relationships between events Wenzel and Jatowt (2023a). (Zhou et al., 2019) categorises temporal commonsense reasoning with respect to five aspects (duration, temporal ordering, typical time, frequency and stationarity). Subsequently, Tan et al. (2023a); Jain et al. (2023) explore the temporal reasoning capabilities of Large Language Models (LLMs) with respect to the temporal commonsense aspects. LLMs with a strong understanding of temporal context can perform better on downstream tasks, including storytelling, natural language inference, timeline comprehension and tracking user status Jain et al. (2023). Thus temporal commonsense reasoning is beneficial for timeline summarisation, as it helps maintain temporal consistency and the correct event order Wenzel and Jatowt (2023b); Vashishtha et al. (2020). Recent work Chan et al. (2024); Feng et al. (2023); Chen et al. (2024); Zhang et al. (2024) has primarily focused on improving the temporal reasoning capabilities of LLMs, without exploring how enhancing these abilities impacts downstream tasks, such as timeline summarisation.
Here we propose combining temporal reasoningwith timeline summarisation using LLMs, to enhance the generation of timeline summaries. Specifically, we first fine-tune a teacher model using a novel temporal reasoning dataset (NarrativeReason) and then distill temporal reasoning knowledge into a smaller student model, which is simultaneously fine-tuned on the timeline summarisation task.
We make the following contributions:
  • $\bullet$
    We are the first to explore how enhancing temporal reasoning in LLMs can improve timeline summarisation.
  • $\bullet$
    Based on the timelines created by NarrativeTime on the TimeBankNT corpus Rogers et al. (2024), we develop a new dataset NarrativeReason for temporal reasoning. Unlike existing temporal reasoning datasets Tan et al. (2023a); Chu et al. (2024); Wang and Zhao (2024), NarrativeReason focuses on the temporal relationships among a series of events within a story rather than event pairs. This can help LLMs process a series of events to generate a coherent and accurate timeline summary.
  • $\bullet$
    We apply the fine-tuned LLM to a completely different domain from what it is trained on, specifically generation of mental health related timeline summaries. Experimental results show that our model achieves the best performance on the timeline summarisation dataset by Song et al. (2024). Not only does it generate more accurate summaries but it also reduces hallucinations in LLMs.
  • $\bullet$
    We show why knowledge distillation works well, and how it induces better learned representations, through activation analysis of the fine-tuned LLM.
2 Related Work
Temporal reasoning for LLMs Temporal reasoning in Natural Language Processing (NLP) is the ability to understand and process information related to time within natural language text. It includes reasoning about the chronology and duration of events, and understanding and capturing different temporal relations Vashishtha et al. (2020). Despite the impressive performace of Large Language Models (LLMs), like GPT-4, across a wide range of tasks (e.g. translation, generation), they have been shown to perform suboptimally in temporal reasoning Wang and Zhao (2024); Chu et al. (2024); Qiu et al. (2023). However the ability to perform temporal reasoning is crucial for understanding narratives Nakhimovsky (1987); Jung et al. (2011a); Cheng et al. (2013), answering questions Bruce (1972); Khashabi (2019); Ning et al. (2020), and summarising events Jung et al. (2011b); Vashishtha et al. (2020). Consequently, efforts are being made to enhance the temporal reasoning capabilities of LLMs. Xing and Tsang (2023) Huang et al. (2024). To increase understanding of temporal expressions, Tan et al. introduced the TEMPREASON dataset which addresses three types of relations (time-time, time-event, event-event). TEMPREASON was used to fine-tune a LLM to improve the model’s temporal reasoning, and they analysed the performance of different LLMs on this dataset showing it is challenging for LLMs to capture the temporal relationships between different events. Xiong et al. use the aligned timeline to help the LLM to understand temporal reasoning by translating the context into a temporal graph,identifying valid time expressions and generating related temporal knowledge. The temporal relationship between events is inferred based on specific times (e.g., the year the events occurred). However in a narrative, events often occur without a clear indication of time.
Temporal reasoning for summarization Jung et al. developed a natural language understanding (NLU) system with a temporal reasoning component to create comprehensive timelines, applied tomedical records, presenting medical history in a more intuitive way. They found that temporal reasoning in NLU is tightly integrated into the NLP system’s deep semantic analysis. Temporal reasoning can help the LM analyze temporal relationships between different events, which is very beneficial for event or news summarisation Vashishtha et al. (2020). However, few studies explore how improvements in temporal reasoning in LLMs directly benefit downstream tasks such as text summarisation.
3 Methodology
Task Given an individual’s timeline (a series of posts between two dates Tsakalidis et al. (2022)), the goal is to generate an abstractive summary that reflects changes in the individual over time Song et al. (2024).
Figure 1: Overview of proposed method. (a) represents fine-tuning the teacher model on the temporal reasoning dataset.
3.1 Proposed architecture
To generate timeline summaries on social media we consider two sub-processes (see Fig. 1):
(1) Improving temporal reasoning. We fine-tune a large LLM as a teacher model on the ‘NarrativeReason’ dataset §3.2.
(2) After fine-tuning the teacher model, we freeze its parameters. At this stage, we fine-tune a student model (a smaller LLM) on the the task of timeline summarisation Chen et al. (2023a) from news. During this process, the teacher model transfers temporal reasoning knowledge to the student model, while the student simultaneously leverages the acquired temporal reasoning knowledge to perform timeline summarisation. For knowledge distillation (KD), we adopt three different strategies: Neuron Selectivity Transfer (NST), Contrastive Representation Distillation (CRD) and Probabilistic Knowledge Transfer (PRT). When generating the timeline summary, we conduct experiments on the TalkLife dataset, a very different domain to the one the student is trained on. We prompt the student model to generate mental health related summaries pertaining to aspects such as diagnostic states, inter- and intra- personal relationships and fluctuations in mood Song et al. (2024).
3.2 Teacher Model
Here the goal is to improve an LLM’s temporal reasoning. There is evidence showing that fine-tuning on datasets such as TEMPLAMA may enable an LLM to memorise the most frequent answer rather develop temporal reasoning Tan et al. (2023b). In other words, the model does not truly understand temporal relationships, such as "before" and "after". Most temporal reasoning datasets involve pairs of events rather than multiple events. However, processing a sequence of events requires more intricate reasoning, including recognising patterns, dependencies, and causal chains among multiple events. This is useful for more sophisticated tasks such as narrative comprehension and timeline summarisation, where understanding the full sequence of events is crucial.
To prevent the LLM from learning shortcuts and memorising the most frequent answer, we created a temporal reasoning dataset NarrativeReason, which contains relationships between a series of events based on a given story.
Event extraction To create NarrativeReason we restructured the NarrativeTime dataset Rogers et al. (2024), which re-annotated TimeBankDense Cassidy et al. (2014) with a timeline-based annotation framework, NarrativeTime. They annotated all possible temporal links (TLINKS) and the temporal relationships of all events occurring within astory. They thus provide a complete temporal relation for the sequence of events in the text, rather than just the temporal relation between event pairs. That means they provide a clear timeline for these events. However, they only used one word (e.g. "fallen") to represent an event, which does not fully capture the complete picture of the event (e.g. "the Indonesian stock market has fallen by twelve percent"). In general, an event is considered an action involving its corresponding participants, and is marked by annotating a representative expression of the event. In the context of temporal reasoning, events are usually the head of the verb phrase. Ning et al. (2018); Pustejovsky et al. (2003).
Therefore, we filtered the dataset TimeBankNT, keeping only annotated verbs to represent events. In order To represent a complete event, we use these verbs as triggers to extract relational triplets e.g. <Indonesian stock market value, fallen, by twelve percent> to represent the event fallen. Then, we use these triplets to construct the temporal relationship of a series event. (e.g. Event <Indonesian stock market value, fallen, by twelve percent> is BEFORE Event <financial week, turning, bad for Asia>).
Dataset construction In a story, we consider the temporal relations for all events, and construct event-event relation question/answer pairs, which addrees the chronological relation between events, such as ‘before’, ‘after’, ‘during’, and ‘simultaneous’ Tan et al. (2023a). Specifically, we obtain the temporal relations of all events and then use questioanswering prompts to reconstruct the dataset. Question: your task is to identify the temporal relation between EVENT A and EVENT B: based on the Story: STORY. Answer: EVENT A temporal relation (BEFORE/ AFTER/ INCLUDES/ IS_INCLUDED/ SIMULTANEOUS) EVENT B. Although a single question-answer pair is used to determine the temporal relationship between a pair of events, for a complete story, we construct multiple question-answer pairs to cover the temporal relationsamong all events. This ensures that the model can clearly outline the temporal relations between all events in the story.
Fine-tuning task We apply supervised fine-tuning (SFT) on a large LLM (teacher model) utilising Low-Rank Adaptation (LoRA) Hu et al. (2022). The input and output of the model are the temporal questions and corresponding answers. Our experiments reveal that fine-tuning on this dataset brings improvements to other temporal reasoning tasks (Appendix A.2) .
Figure 2: The temporal relationship between events. We represent events triggered by verbs using relational triplets, as shown in the middle column. In the right column, we list a series of events and use arrows of different colors to represent the temporal relations between these events. In this example, we use blue, green and purple arrows to indicate ’Simultaneous’, ’After’ and ’Before’ respectively.
3.3 Student Model
After we fine-tune a teacher model on the reconstructed dataset, we transfer the temporal reasoning knowledge to a student model, while, also fine-tuning the student on a news timeline summarisation dataset Chen et al. (2023b). Thus we aim for the student to learn temporal reasoning and use this ability to generate timeline summaries. We fine-tune Phi-3-mini-4k-instruct as a student model. We use three knowledge distillation (KD) objectives to transfer knowledge from the teacher to the student: Neuron Selectivity Transfer (NST) Huang and Wang (2017), transfers heatmap like spatial activation patterns of teacher neurons to student neurons; Contrastive Representation Distillation (CRD) Tian et al. (2019), maximises the mutual information between the teacher and student representations with contrastive learning; Probabilistic Knowledge Transfer (PRT) Passalis and Tefas (2018), matches the probability distribution of the data in the feature space.
PRT:
The core idea is to match the probability distribution of the data in the feature space between teacher and student models. However, learning a significantly smaller model that accurately recreates the whole geometry of a complex teacher model is often impossible. Passalis and Tefas (2018) uses the conditional probability distribution to describe the samples. Here, $\mathbf{Y_{t}}={\{\mathbf{y_{t}}}_{1},{\mathbf{y_{t}}}_{2},...,{\mathbf{y_{t}}}_{l}\}\in\mathbb{R}^{vocab_{t}}$ denote the output logits of the teacher model, and $\mathbf{Y_{s}}={\{\mathbf{y_{s}}}_{1},{\mathbf{y_{s}}}_{2},...,{\mathbf{y_{s}}}_{l}\}\in\mathbb{R}^{vocab_{s}}$ denote the output logits of the student model, where ${\mathbf{y_{t}}}$/${\mathbf{y_{s}}}$ is vector and $l$ is the length of sentences, ${vocab_{t}}$ and ${vocab_{t}}$ are the vocabulary sizes of teacher and student models respectively. We can define the conditional probability distribution for the teacher model as Eq. 1, and student model as Eq. 2.
${p_{i|j}}=\frac{K({\mathbf{y_{t}}_{i}},\mathbf{y_{t}}_{j};2\sigma_{t}^{2})}{\sum_{k=1,k\neq{j}}^{l}{K(\mathbf{y_{t}}_{k},\mathbf{y_{t}}_{j};2\sigma_{t}^{2})}}$ (1)
${q_{i|j}}=\frac{K(\mathbf{x_{s}}_{i},\mathbf{x_{s}}_{j};2\sigma_{s}^{2})}{\sum_{k=1,k\neq{j}}^{l}{K(\mathbf{x_{s}}_{k},\mathbf{x_{s}}_{j};2\sigma_{s}^{2})}}$ (2)
In these two equations (Eq. 1 and Eq. 2), we use the cosine similarity as a kernel metric allows for more robust affinity estimations.
${K_{consine}({\mathbf{y_{t}}_{i}},{\mathbf{y_{t}}_{j}})}=\frac{1}{2}(\frac{{\mathbf{y_{t}}_{i}^{\mathrm{T}}}\mathbf{y_{t}}_{j}}{\|{\mathbf{y_{t}}_{i}}\|{{}_{2}}\|{\mathbf{y_{t}}_{j}}\|{{}_{2}}}+1)\in[0,1].$
We use Kullback-Leibler (KL) divergence to calculate the distance between the conditional probability distributions of the teacher model and the student model, as follows:
${L_{PKT}}={\sum_{i=1}^{l}\sum_{j=1,i\neq{j}}^{l}}{p_{i|j}}\log({\frac{p_{i|j}}{q_{i|j}}}).$
NST
matches the distributions of neuron selectivity patterns between teacher and student networks. We transfer the last hidden layer $\mathbf{T}=\mathbf{t(x)}$ of the teacher model to the last hidden layer $\mathbf{S}=\mathbf{s(x)}$ of the student model given input text $\mathbf{x}$. Specifically, we transfer neuron selectivity knowledge from $\{{\mathbf{t(x)}}_{*,i}\}_{i=1}^{N}$ to $\{{\mathbf{s(x)}}_{*,i}\}_{i=1}^{M}$, where $N$ and $M$ are the hidden state dimensions. Then we follow the idea in Huang and Wang (2017) using Maximum Mean Discrepancy (MMD) to calculate the distance between these two distributions. The basic idea of MMD is that if the moments of two random variables are the same for all orders, then the two distributions are identical $L_{MMD^{2}}(\mathbf{t},\mathbf{s})$.
${L_{MMD^{2}}(\mathbf{t},\mathbf{s})}={\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{i^{\prime}=1}^{N}k\left[{\mathbf{t(x)}}_{*,i};{\mathbf{t(x)}}_{*,i^{\prime}}\right]}$
$+{{\frac{1}{M^{2}}\sum_{j=1}^{M}\sum_{j^{\prime}=1}^{M}k\left[{\mathbf{s(x)}}_{*,j};{\mathbf{s(x)}}_{*,j^{\prime}}\right]}}$
$-{{\frac{1}{MN}\sum_{i=1}^{N}\sum_{j=1}^{M}k\left[{\mathbf{s(x)}}_{*,i};{\mathbf{t(x)}}_{*,j}\right]}}.$
In this formula we use a Gaussian Kernel ${k(x,y)=exp(-\frac{{||x-y||}_{2}^{2}}{2{\sigma}^{2}})}$ with $\sigma=1$. We transfer the teacher activation patterns to the student by minimizing $L_{MMD^{2}}$.
CRD
maximises the lower-bound to the mutual information between the teacher and student representations. Here, we follow Tang et al. (2021) to sample one positive pair from the joint distribution $p(\mathbf{S},\mathbf{T})=q(\mathbf{S},\mathbf{T}|positive)$ for $N$ negative pairs sampled from the product of marginal $p(\mathbf{S})p(\mathbf{T})=q(\mathbf{S},\mathbf{T}|negtive)$, where $N$ is the batch size. We can maximize the lower bound of mutual information by minimizing the following loss function:
${L_{CRD}(\mathbf{x})}=-\mathbb{E}_{q(\mathbf{s},\mathbf{t}|positive)}\left[\log h(\mathbf{s},\mathbf{t})\right]$ (3)
$-N.\mathbb{E}_{q(\mathbf{s},\mathbf{t}|negtive)}\left[\log(1-h(\mathbf{s},\mathbf{t}))\right]$
In Eq 3, function $h$ should satisfy $h:\{\mathbf{s},\mathbf{t}\}\rightarrow[0,1]$,
${h(\mathbf{s},\mathbf{t})}=\frac{exp({\mathbf{s}^{\mathrm{T}}}\mathbf{t})}{exp({\mathbf{s}^{\mathrm{T}}}\mathbf{t})+\frac{N}{M}},$
where M is the cardinality of the dataset, and we need to normalize $\mathbf{s}$ and $\mathbf{t}$ by $L$-2 norm before the inner product.
The knowledge distillation method transfers temporal reasoning knowledge from the teacher model to the student model. At the same time, we aim for this knowledge to benefit the timeline summarization task. Therefore, we fine-tune the student model on the timeline summarization dataset §4.1, enabling it to simultaneously learn from the teacher model and use the language modeling loss (for next token prediction) $L_{language}$ to integrate temporal reasoning knowledge with timeline summarization information.
3.4 Mental Health Timeline Summary
We apply the student model to other domains, specifically to generate mental health-related summaries for timelines §4.1 from social media. For the mental health summary, we use the format proposed by Song et al. (2024), which includes three key clinical concepts (diagnosis, inter- and intra- personal relations, moments of change). We follow their method to prompt the student model to generate a summary for each timeline.
4 Experiments
4.1 Datasets
We conduct experiments on three different datasets. We fine-tune the teacher model on the ‘NarrativeReason’ dataset. When distilling the temporal reason knowledge to the student model, we use a news timeline summarisation datasetChen et al. (2023a). Finally, we apply the model to a different domain, specifically generating mental health-related timeline summaries from social media.
NarrativeReason We extracted 668 events from 30 articles, containing a total of 19,614 temporal relations between events Rogers et al. (2024), leading to 19,614 question/answer pairs for event-event relations. We use these question/answer pairs to fine-tune the teacher model to enhance its temporal reasoning capability.
Timeline summarisation Dataset This timeline summarisation dataset is sourced from Chen et al. (2023a). They collected timeline summaries from Wikipedia websites, with a total of 5,000 timelines and summaries.
TalkLife When generating the summary, we use the dataset collected by Tsakalidis et al. (2022) comprising 500 anonymised user timelines from Talklife. Song et al. (2024) sample 30 timelines from it and annotate them with corresponding mental health-related summaries. These human-written summaries include diagnosis, intra- and interpersonal patterns and mental state changes over time, and they also highlight information related to individuals’ mental states, which can be used for automated evaluation.
4.2 Models & Baselines
We compare our method against existing LLMs for mental health related summarisation. We introduce the implementation details in the appendix A.1
L-phi This model uses LLaMA as the teacher model and a smaller model, Phi, as the student. We apply different knowledge distillation (KD) methods to transfer temporal reasoning knowledge to the student model. Subsequently, we directly prompt this model to generate mental health-related timeline summaries.
P-phi In this model, we use Phi as the teacher model and another Phi of the same size as the student model.
$\text{Phi}_{joint}$ To compare with the phi-phi, we use joint learning to fine-tune Phi on both the NarrativeReason and Timeline summarisation datasets. This allows us to observe whether knowledge distillation (KD) outperforms directly fine-tuning on the two datasets. In addition, we fine-tune Phi on the NarrativeReason and timeline summarisation datasets and obtain models $\text{Phi}_{temp}$ and $\text{Phi}_{timeline}$ separately. We use these two models to generate timeline summaries for comparison with the KD derived model.
$\text{Phi}_{ICL}$ We use in context learning (ICL) to guide Phi to generate summaries. We provide the model with a pair consisting of a timeline and its corresponding summary as an example, and then let it generate summaries for other timelines.
LLaMA We prompt LLaMA to generate mental health related summaries on timelines directly as a baseline.
4.3 Evaluation
We use the timeline summaries from Song et al. (2024)4.1) forautomated evaluation. We employFactual Consistency (FC), to measure whether timeline summaries are consistent with the original timelines, and Evidence Appropriateness (EA), to measure the consistency of human written summaries with their corresponding timeline summaries Song et al. (2024).
For human evaluation, we worked with two clinical psychology graduate students fluent in English to evaluate 30 summaries generated from 30 timelines (TalkLfe). We follow the metrics used in Song et al. (2024) to evaluate the summaries from the perspectives of Factual Consistency and Usefulness (general/diagnosis/inter-&Interpersonal/MOC). A factually consistent summary should accurately represent the content of the timeline and excludes any information not present in the timeline. We also evaluate the usefulness of a mental health summary on the basis of four aspects: General (contains the most clinically important information from the timeline and aids the clinician in understanding a patient’s condition); Diagnosis (provides useful information about an individual’s diagnosis); inter-&Interpersonal (provides helpful information about an individuals’ main needs and patterns of self and other relationships); MOC (provides useful information about an individual’s changes over time in terms of emotion/cognition and behaviour).
5 Results and Discussion
5.1 Automatic evaluation
We conducted experiments with different combinations of KD methods. Table 1 shows experiment results. Since we did not change the output dimensions when fine-tuning LLaMA, the CRD method was not used during the KD process. Among the individual methods, PKT performed the best; however, combining PKT with NST achieved the best overall results.
[…]
Table 1: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) on student model.
Higher is better, best in bold
Table 2 shows the results of applying alternative fine-tuning strategies on LLMs. The results for $\text{Phi}_{temp}$ and $\text{Phi}_{timeline}$ on FC indicate that fine-tuning on a single dataset does not improve model performance; instead, it exacerbates hallucination issues. Notably, $\text{Phi}_{temp}$ performed the worst in both FC and EA metrics, suggesting that the incorporation of temporal reasoning information appears to interfere with the LLM’s ability to effectively handle the timeline summarization task. Additionally, in strategy $\text{Phi}_{jont}$, combining the two types of data directly during training failed to integrate them effectively. As a result, the performance of $\text{Phi}_{jont}$ was even worse than using in-context learning to guide the LLM ($\text{Phi}_{ICL}$).
Metric $\text{Phi}_{ICL}$ $\text{Phi}_{temp}$ $\text{Phi}_{timeline}$ $\text{Phi}_{joint}$ FC .412 .141 .184 .238 EA .965 .895 .966 .941
Table 2: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) across the compared models. Higher is better, best in bold
5.2 Human evaluation
Based on the results of the automatic evaluation, we selected the best-performing LLaMA-Phi and Phi-Phi models. Additionally, we included the non-fine-tuned versions of LLaMA and Phi. This can help us understand in which specific aspects the model has improved with the inclusion of temporal reasoning information. The fine-tuned model shows the greatest improvement in terms of factual consistency and usefulness (general). This aligns with our findings when analyzing the summaries, where the fine-tuned model significantly reduces hallucination, as shown in Table 3. In addition, we found that the fine-tuned model did not show significant improvement in terms of Moments of Change (MOC). This could be because, in the timelines, users do not experience many emotional switches, resulting in the model-generated summaries being relatively similar.
Aspect phi phi-phi LLaMA LLaMA-phi Factual Consistency 2.90 3.32 3.58 3.83 Usefulness (General) 2.60 3.13 3.17 3.48    (Diagnosis) 2.90 3.37 3.45 3.62   —-(Inter-& Intrapersonal) 2.95 3.00 3.40 3.51   —-(MoC) 2.97 2.97 3.42 3.47
Table 3: Human evaluation results based on 5-point Likert scales (1 is worst, 5 is best). Best in bold.
5.3 Why knowledge distillation works
In this section, we analyze why the $\text{Phi}_{timeline}$ model performed better from a representation learning perspective. For this purpose, we perform two experiments, a task understanding probing experiment and a Joint Task Representation Learning (JTRL) experiment. We analyze the internal representations of $\text{Phi}_{timeline}$ against $\text{Phi}_{joint}$. We construct our probing dataset by using $\text{Phi}_{joint}$’s test dataset. We pose the two tasks as a binary classification problem and then extract the activations for the last layer. Then we use UMAP (McInnes et al., 2018) to project these activations to lower dimensions (Sainburg et al., 2021; Tseriotou et al., 2023). From Figure 3, we can see that the activations of $\text{Phi}_{joint}$ are well separated for each task, whereas the activations for $\text{Phi}_{timeline}$ overlap. Given the performance of $\text{Phi}_{timeline}$, this leads us to hypothesize that the model learned better representations for the task due to more task specific polysemantic (Olah et al., 2020) neurons.
To validate our hypothesis, we ran another set of experiments to analyze the internal representation difference between the two models, which we termed as JTRL. Our main idea here was to validate that knowledge distillation resulted in better representations due to more polysemantic neurons, and these representations varied highly compared to $\text{Phi}_{joint}$’s representation. To measure JTRL, we use the Centered Kernel Alignment (CKA) (Kornblith et al., 2019) similarity score. CKA can be used to measure the similarity between internal representations of modelsConneau et al. (2020); Muller et al. (2021); Del and Fishel (2021); Moosa et al. (2023).
To calculate CKA, we used the same probing dataset. First, we calculate the sentence embeddings of each input by averaging the hidden state representation of the tokens. Then, we calculate the CKA similarity score between the mean sentence embeddings and each layer representation of the model. We did this both for the individual models and between the layers of the models. From Figure 4, we can see that $\text{Phi}_{timeline}$ shows a gradual increase in CKA across layers, peaking in the mid-to-late layers. This indicates that as the layers progress, $\text{Phi}_{timeline}$ preserves and refines the information from the initial embeddings in a task-relevant way. The is likely due to the fact that the distillation process encourages this alignment by transferring task-relevant knowledge from the teacher. On the other hand, $\text{Phi}_{joint}$ shows an initial increase but saturates and flattens early. This indicates that it fails to refine representations effectively in deeper layers, which could be due to conflicting objectives between the tasks. The lower CKA in later layers suggests that the model moves away from the initial embeddings in a manner that is less effective for task-specific learning. Lastly, the CKA values between the models show that they learn vastly different representations. In short, $\text{Phi}_{timeline}$’s ability to align with the initial embeddings correlates with its better task performance, as the CKA value reflects how well the model retains and transforms meaningful input information throughout its layers.
(a) UMAP projection $\text{Phi}_{joint}$
(b) UMAP projection $\text{Phi}_{timeline}$
[…]
Figure 4: CKA similarity score of both within and between $\text{Phi}_{timeline}$ and $\text{Phi}_{joint}$ model representations.
6 Conclusions
We created a dataset named NarrativeReason, which is utilised to enhance the temporal reasoning abilities of LLMs. This dataset enables models to uncover temporal relationships between events within timelines. We fine-tune a larger LLM with NarrativeReason, then distill its enhanced temporal reasoning capability to a smaller LLM while leveraging the distilled knowledge to enhance performance in timeline summarisation tasks. We apply the model to a different domain from the one it was trained on, namely generating mental health related timeline summaries. Our results demonstrate that our approach using knowledge distillation produces more accurate summaries while significantly reducing hallucinations. We further analyzed the internal representations of the model and found that knowledge distillation leads to better feature representations within the model that are more aligned with the task of timeline summarisation.
Limitations
In our work we aim to leverage temporal reasoning to enhance the performance of LLMs in timeline summarisation tasks. We apply this to social media timelines from the mental health domain to enable LLMs to recognize and analyze events chronologically, in order to capture the dynamics of user behavior and evolving mental states more effectively. This faces the following limiting factors: (a) the existing knowledge of mental health embedded in the LLM and (b) the fixed formats of mental health-related texts that the model is trained on. By analyzing the generated summaries, we found that there seems to be a general tendency to make clear statements about specific DSM diagnoses (such as PTSD, bipolar disorder, etc.). In the vast majority of cases it is possible to write that there is evidence that can indicate such a potential, instead of providing a definite assessment. Moreover, many parts of the summaries seem very generic. This lack of personalisation can sometimes lower the quality of the summary. While this may not always have a significant impact, it generally reduces the depth and individuality of the analysis, sometimes even affecting the factual consistency.
These findings can help the exploration of LLMs in the future, particularly in the mental health domain. They can help refine future models that better understand social media posts, leading to more accurate mental health summaries and improved diagnostic insights for clinicians.
Ethics Statement
Ethics institutional review board (IRB) approval was obtained from the corresponding ethics board of the lead University prior to engaging in this research study. Our work involves ethical considerations around the analysis of user generated content shared on a peer support network (TalkLife). A license was obtained to work with the user data from TalkLife and a project proposal was submitted to them in order to embark on the project.
The final summaries in all cases are obtained by feeding the timeline summaries into an LLM. Given that LLMs are susceptible to factual inaccuracies, often referred to as ’hallucinations,’ and tend to exhibit biases, the clinical summaries they generate may contain errors that could have serious consequences in the realm of mental health decision-making. These inaccuracies can encompass anything from flawed interpretations of the timeline data to incorrect diagnoses and even recommendations for potentially harmful treatments. Mental health professionals must exercise caution when relying on such generated clinical summaries. These summaries should not serve as substitutes for therapists in making clinical judgments. Instead, well-trained therapists must skillfully incorporate these summaries into their clinical thought processes and practices. Significant efforts are required to establish the scientific validity of the clinical benefits offered by these summaries before they can be integrated into routine clinical practice.
Appendix A Appendix
A.1 Implementation Details
We use Meta-Llama-3-8B as teacher model and Phi-3-mini-4k-instruct as student model. We use Low-Rank Adaptation (LoRA) Hu et al. (2022) for both of these two models while fine-tuning. We use an AdamW Kingma and Ba (2015) optimizer with learning rate 5e-5. While fine-tuning, we set the batch-size as 1 for each task, but set gradient accumulation steps as 16.
A.2 Experiment on teacher model
After fine-tuning the teacher model, we used the TEMPREASON dataset Tan et al. (2023b) to evaluate its performance pre- and post-fine-tuning. Specifically, we focused on the L3 part of the dataset, which deals with event-event relations, to determine whether the model could accurately infer the temporal sequence between two events. We use F1 as the evaluation metrics, the experimental results show that the fine-tuned model achieved a 0.07 improvement in this metric compared to the pre-fine-tuning model, highlighting its enhanced ability to infer event-event temporal relations.

### v2

@@
 Abstract
-This paper explores whether enhancing temporal reasoning capabilities in Large Language Models (LLMs) can improve the quality of timeline summarization, the task of summarising long texts containing sequences of events, particularly social media threads .
-We introduce NarrativeReason, a novel dataset focused on temporal relationships among sequential events within narratives, distinguishing it from existing temporal reasoning datasets that primarily address pair-wise event relationships.
-Our approach then combines temporal reasoning with timeline summarization through a knowledge distillation framework, where we first fine-tune a teacher model on temporal reasoning tasks and then distill this knowledge into a student model while simultaneously training it for the task of timeline summarization.
-Experimental results demonstrate that our model achieves superior performance on mental health-related timeline summarization tasks, which involve long social media threads with repetitions of events and a mix of emotions, highlighting the importance of leveraging temporal reasoning to improve timeline summarisation.
+This paper explores whether enhancing temporal reasoning capabilities in Large Language Models (LLMs) can improve the quality of timeline summarisation, the task of summarising long texts containing sequences of events, such as social media threads.
+We first introduce NarrativeReason, a novel dataset focused on temporal relationships among sequential events within narratives, distinguishing it from existing temporal reasoning datasets that primarily address pair-wise event relationships.
+Our approach then combines temporal reasoning with timeline summarisation through a knowledge distillation framework, where we first fine-tune a teacher model on temporal reasoning tasks and then distill this knowledge into a student model while simultaneously training it for the task of timeline summarisation.
+Experimental results demonstrate that our model achieves superior performance on out-of-domain mental health-related timeline summarisation tasks, which involve long social media threads with repetitions of events and a mix of emotions, highlighting the importance and generalisability of leveraging temporal reasoning to improve timeline summarisation.
 1 Introduction
-Timeline summarization organizes and presents a sequence of events in a coherent and concise manner Steen and Markert (2019); Li et al.
+Timeline summarisation organizes and presents a sequence of events in a coherent and concise manner Steen and Markert (2019); Li et al.
 (2021); Hu et al.
@@
 Temporal reasoning involves understanding and processing temporal information in text to deduce time-based relationships between events Wenzel and Jatowt (2023a).
-(Zhou et al., 2019) categorises temporal commonsense reasoning with respect to five aspects (duration, temporal ordering, typical time, frequency and stationarity).
+Zhou et al.
+(2019) categorises temporal commonsense reasoning with respect to five aspects (duration, temporal ordering, typical time, frequency and stationarity).
 Subsequently, Tan et al.
-(2023a); Jain et al.
-(2023) explore the temporal reasoning capabilities of Large Language Models (LLMs) with respect to the temporal commonsense aspects.
+(2023); Jain et al.
+(2023) explore the temporal reasoning capabilities of Large Language Models (LLMs) with respect to temporal commonsense aspects.
 LLMs with a strong understanding of temporal context can perform better on downstream tasks, including storytelling, natural language inference, timeline comprehension and tracking user status Jain et al.
@@
 (2020).
-Recent work Chan et al.
+Despite the evidenced connection between temporal reasoning and timeline summarisation, recent work Chan et al.
 (2024); Feng et al.
@@
 (2024); Zhang et al.
-(2024) has primarily focused on improving the temporal reasoning capabilities of LLMs, without exploring how enhancing these abilities impacts downstream tasks, such as timeline summarisation.
+(2024) has primarily focused on improving the temporal reasoning capabilities of LLMs, without exploring how they impact downstream tasks, such as timeline summarisation.
 Here we propose combining temporal reasoning with timeline summarisation using LLMs, to enhance the generation of timeline summaries.
-Specifically, we first fine-tune a teacher model using a novel temporal reasoning dataset (NarrativeReason) and then distill temporal reasoning knowledge into a smaller student model, which is simultaneously fine-tuned on the timeline summarisation task.
+Specifically, we first fine-tune a teacher model using a novel temporal reasoning dataset (NarrativeReason) and then distill temporal reasoning knowledge into a smaller student model, which is simultaneously fine-tuned on the timeline summarisation task, trained and tested in separate domains.
 We make the following contributions:
@@
   • $\bullet$
-    Based on the timelines created by NarrativeTime on the TimeBankNT corpus Rogers et al.
-(2024), we develop a new dataset NarrativeReason for temporal reasoning.
+    Based on the timelines derived in NarrativeTime Rogers et al.
+(2024) from the TimeBankNT corpus Cassidy et al.
+(2014), we develop a new dataset for temporal reasoning, NarrativeReason.
 Unlike existing temporal reasoning datasets Tan et al.
-(2023a); Chu et al.
-(2024); Wang and Zhao (2024), NarrativeReason focuses on the temporal relationships among a series of events within a story rather than event pairs.
+(2023); Chu et al.
+(2024); Wang and Zhao (2024), NarrativeReason focuses on the temporal relationships among a series of events within a story rather than distinct event pairs.
 This can help LLMs process a series of events to generate a coherent and accurate timeline summary.
   • $\bullet$
-    We apply the fine-tuned LLM to a completely different domain from what it is trained on, specifically generation of mental health related timeline summaries.
+    We fine-tune a large LLM on NarrativeReason and distill its knowledge to a smaller model, which is fine-tuned for the task of timeline summarisation.
+The resulting fine-tuned smaller LLM is applied to a completely different domain from the one it is trained on.
 Experimental results show that our model achieves the best performance on the timeline summarisation dataset by Song et al.
@@
   • $\bullet$
-    We show why knowledge distillation works well, and how it induces better learned representations, through activation analysis of the fine-tuned LLM.
+    We show why knowledge distillation works well, and how it induces better learned representations, through activation analysis of the fine-tuned smaller LLM.
 2 Related Work
@@
 (2020).
-Despite the impressive performace of Large Language Models (LLMs), like GPT-4, across a wide range of tasks (e.g.
-translation, generation), they have been shown to perform suboptimally in temporal reasoning Wang and Zhao (2024); Chu et al.
+Despite the impressive performance of Large Language Models (LLMs), like GPT-4, across a wide range of tasks (e.g.
+translation, generation), they have been shown to perform sub-optimally in temporal reasoning Wang and Zhao (2024); Chu et al.
 (2024); Qiu et al.
@@
 (2020).
-Consequently, efforts are being made to enhance the temporal reasoning capabilities of LLMs.
-Xing and Tsang (2023) Huang et al.
+Consequently, efforts are being made to enhance the temporal reasoning capabilities of LLMs Xing and Tsang (2023) Huang et al.
 (2024).
 To increase understanding of temporal expressions, Tan et al.
-introduced the TEMPREASON dataset which addresses three types of relations (time-time, time-event, event-event).
-TEMPREASON was used to fine-tune a LLM to improve the model’s temporal reasoning, and they analysed the performance of different LLMs on this dataset showing it is challenging for LLMs to capture the temporal relationships between different events.
+(2023) introduced the TEMPREASON dataset which addresses three types of relations (time-time, time-event, event-event).
+TEMPREASON was used to fine-tune a LLM to improve its temporal reasoning, and performance of different LLMs on this dataset showed it is challenging for LLMs to capture the temporal relations between different events.
 Xiong et al.
-use the aligned timeline to help the LLM to understand temporal reasoning by translating the context into a temporal graph,identifying valid time expressions and generating related temporal knowledge.
-The temporal relationship between events is inferred based on specific times (e.g., the year the events occurred).
+(2024) use an aligned timeline to improve an LLM’s temporal reasoning by translating the context into a temporal graph,identifying valid time expressions and generating related temporal knowledge.
+The temporal relationship between events is inferred based on specific times (e.g., year of event).
 However in a narrative, events often occur without a clear indication of time.
-Temporal reasoning for summarization Jung et al.
-developed a natural language understanding (NLU) system with a temporal reasoning component to create comprehensive timelines, applied tomedical records, presenting medical history in a more intuitive way.
-They found that temporal reasoning in NLU is tightly integrated into the NLP system’s deep semantic analysis.
-Temporal reasoning can help the LM analyze temporal relationships between different events, which is very beneficial for event or news summarisation Vashishtha et al.
+Temporal reasoning for summarisation Jung et al.
+(2011b) developed a natural language understanding (NLU) system with a temporal reasoning component to create comprehensive timelines, applied to medical records, presenting medical history in a more intuitive way.
+They found that temporal reasoning in NLU is tightly integrated into the NLP system’s deep semantic analysis and can help a LM analyze temporal relationships between different events, which is beneficial for event or news summarisation Vashishtha et al.
 (2020).
-However, few studies explore how improvements in temporal reasoning in LLMs directly benefit downstream tasks such as text summarisation.
+However, despite the evidenced connection, few studies explore how improvements in temporal reasoning in LLMs directly benefit downstream tasks such as text summarisation.
 3 Methodology
@@
 (2024).
-Figure 1: Overview of proposed method. (a) represents fine-tuning the teacher model on the temporal reasoning dataset.
+Figure 1: Overview of proposed method.
+(a) represents fine-tuning the teacher model on the temporal reasoning dataset; (b) In the KD process, we input the timeline summarisation dataset into both the Teacher and Student models, transferring temporal reasoning knowledge while using it to assist the timeline summarisation task to fine-tune Student model.
 3.1 Proposed architecture
 To generate timeline summaries on social media we consider two sub-processes (see Fig. 1):
-(1) Improving temporal reasoning. We fine-tune a large LLM as a teacher model on the ‘NarrativeReason’ dataset §3.2.
+(1) Improving temporal reasoning. We fine-tune a large LLM as a teacher model on our ‘NarrativeReason’ dataset §3.2.
 (2) After fine-tuning the teacher model, we freeze its parameters.
-At this stage, we fine-tune a student model (a smaller LLM) on the the task of timeline summarisation Chen et al.
-(2023a) from news.
+At this stage, we fine-tune a student model (a smaller LLM) on timeline summarisation from news Chen et al.
+(2023).
 During this process, the teacher model transfers temporal reasoning knowledge to the student model, while the student simultaneously leverages the acquired temporal reasoning knowledge to perform timeline summarisation.
 For knowledge distillation (KD), we adopt three different strategies: Neuron Selectivity Transfer (NST), Contrastive Representation Distillation (CRD) and Probabilistic Knowledge Transfer (PRT).
-When generating the timeline summary, we conduct experiments on the TalkLife dataset, a very different domain to the one the student is trained on.
-We prompt the student model to generate mental health related summaries pertaining to aspects such as diagnostic states, inter- and intra- personal relationships and fluctuations in mood Song et al.
+When generating the timeline summary, we conduct experiments on the TalkLife dataset, which pertains to a very different domain (mental health) to the one the student is trained on (news).
+We prompt the student model to generate mental health related summaries pertaining to aspects such as diagnostic states, inter- and intra- personal relationships and fluctuations in mood following Song et al.
 (2024).
-3.2 Teacher Model
+3.2 Teacher Model and NarrativeReason dataset
 Here the goal is to improve an LLM’s temporal reasoning.
-There is evidence showing that fine-tuning on datasets such as TEMPLAMA may enable an LLM to memorise the most frequent answer rather develop temporal reasoning Tan et al.
-(2023b).
-In other words, the model does not truly understand temporal relationships, such as "before" and "after".
-Most temporal reasoning datasets involve pairs of events rather than multiple events.
-However, processing a sequence of events requires more intricate reasoning, including recognising patterns, dependencies, and causal chains among multiple events.
+Evidence has shown that fine-tuning on datasets such as TEMPLAMA may enable an LLM to memorise the most frequent answer rather than develop temporal reasoning Tan et al.
+(2023).
+In other words, the model does not truly learn the meaning of temporal relations, such as "before" and "after".
+We hypothesise that this is because most temporal reasoning datasets involve pairs of events rather than multiple events.
+Processing a sequence of events requires more intricate reasoning, including recognising patterns, dependencies, and causal chains among multiple events.
 This is useful for more sophisticated tasks such as narrative comprehension and timeline summarisation, where understanding the full sequence of events is crucial.
-To prevent the LLM from learning shortcuts and memorising the most frequent answer, we created a temporal reasoning dataset NarrativeReason, which contains relationships between a series of events based on a given story.
+To prevent the LLM from learning shortcuts and memorising the most frequent answer, we created a temporal reasoning dataset NarrativeReason, which contains relations between a series of events based on a given narrative.
 Event extraction To create NarrativeReason we restructured the NarrativeTime dataset Rogers et al.
-(2024), which re-annotated TimeBankDense Cassidy et al.
-(2014) with a timeline-based annotation framework, NarrativeTime.
-They annotated all possible temporal links (TLINKS) and the temporal relationships of all events occurring within astory.
-They thus provide a complete temporal relation for the sequence of events in the text, rather than just the temporal relation between event pairs.
-That means they provide a clear timeline for these events.
-However, they only used one word (e.g.
-"fallen") to represent an event, which does not fully capture the complete picture of the event (e.g.
-"the Indonesian stock market has fallen by twelve percent").
-In general, an event is considered an action involving its corresponding participants, and is marked by annotating a representative expression of the event.
-In the context of temporal reasoning, events are usually the head of the verb phrase.
-Ning et al.
+(2024), which in turn had re-annotated TimeBankDense Cassidy et al.
+(2014) with a timeline-based annotation framework, NarrativeTime framework.
+Rogers et al.
+(2024) had annotated all possible temporal links (TLINKS) between all events occurring within a narrative, thus providing temporal relations for the entire sequence of events (timeline) rather than just between event pairs to get a NarrativeTime dataset.
+Here, TLINKs usually denote the event order information (e.g., before, after, during).
+Events more broadly and especially in the context of temporal reasoning, are represented as relation triples where the event trigger is the head of the verb phrase Ning et al.
 (2018); Pustejovsky et al.
-(2003).
-Therefore, we filtered the dataset TimeBankNT, keeping only annotated verbs to represent events.
-In order To represent a complete event, we use these verbs as triggers to extract relational triplets e.g.
-<Indonesian stock market value, fallen, by twelve percent> to represent the event fallen.
-Then, we use these triplets to construct the temporal relationship of a series event.
+(2003), linking the corresponding arguments.
+However, NarrativeTime dataset  Rogers et al.
+(2024) only denotes the type of the event, annotated at the level of the verb, e.g.
+for the sentence "the value of the Indonesian stock market has fallen by twelve percent" this is annotated in NarrativeTime datasets as "the value of the Indonesian stock market has <EVENT class="OCCURRENCE" eid="e7"> fallen </EVENT> by twelve percent".
+However, this only constitutes a denser temporal annotation without providing event triples and is therefore unsuited for temporal reasoning training tasks.
+Thus we reconstructed NarrativeReason, to augment it with event triples that can be used for temporal reasoning training.
+Specifically, we filtered the NarrativeTime dataset, keeping only annotated verbs to represent events.
+In order to represent an event, we use verbs as triggers to extract relational triples e.g.
+<Indonesian stock market value, fallen, by twelve percent> to represent the event annotated in NarrativeTime dataset as fallen.
+Then, we use these triples to construct the temporal relations of a series event.
 (e.g.
 Event <Indonesian stock market value, fallen, by twelve percent> is BEFORE Event <financial week, turning, bad for Asia>).
-Dataset construction In a story, we consider the temporal relations for all events, and construct event-event relation question/answer pairs, which addrees the chronological relation between events, such as ‘before’, ‘after’, ‘during’, and ‘simultaneous’ Tan et al.
-(2023a).
-Specifically, we obtain the temporal relations of all events and then use questioanswering prompts to reconstruct the dataset.
+Dataset construction For a given narrative, we consider the temporal relations between all events, and construct question/answer pairs for event-event relations, addressing the chronological order of events (‘before’, ‘after’, ‘during’, and ‘simultaneous’ Tan et al.
+(2023)).
+Specifically, we obtain the temporal relations of all events and then use question answering prompts to reconstruct the dataset.
 Question: your task is to identify the temporal relation between EVENT A and EVENT B: based on the Story: STORY.
-Answer: EVENT A temporal relation (BEFORE/ AFTER/ INCLUDES/ IS_INCLUDED/ SIMULTANEOUS) EVENT B.
-Although a single question-answer pair is used to determine the temporal relationship between a pair of events, for a complete story, we construct multiple question-answer pairs to cover the temporal relationsamong all events.
-This ensures that the model can clearly outline the temporal relations between all events in the story.
+Answer: EVENT A temporal relation (BEFORE/ AFTER/ INCLUDES/ IS_INCLUDED/ SIMULTANEOUS) EVENT B Tan et al.
+(2023).
+Although a single question-answer pair is used to determine the temporal relation between a pair of events, for a complete narrative, we construct multiple question-answer pairs to cover the temporal relations among all events.
+This ensures that the model is exposed to all temporal relations between all events in the story.
+Fig.
+2 shows the data format of NarrativeReason.
 Fine-tuning task We apply supervised fine-tuning (SFT) on a large LLM (teacher model) utilising Low-Rank Adaptation (LoRA) Hu et al.
@@
 The input and output of the model are the temporal questions and corresponding answers.
-Our experiments reveal that fine-tuning on this dataset brings improvements to other temporal reasoning tasks (Appendix A.2) .
-Figure 2: The temporal relationship between events.
-We represent events triggered by verbs using relational triplets, as shown in the middle column.
-In the right column, we list a series of events and use arrows of different colors to represent the temporal relations between these events.
-In this example, we use blue, green and purple arrows to indicate ’Simultaneous’, ’After’ and ’Before’ respectively.
+Our experiments show that indeed fine-tuning on NarrativeReason improves performance on established temporal reasoning tasks (Appendix A.2) .
+Figure 2: The temporal relations between events.
+The text in the left column comes from NarrativeTime dataset, and we filtered it to keep only verbs.
+We represent events triggered by verbs using relational triples, as shown in the middle column.
+In the right column, we construct question/answer pairs for all events.
 3.3 Student Model
-After we fine-tune a teacher model on the reconstructed dataset, we transfer the temporal reasoning knowledge to a student model, while, also fine-tuning the student on a news timeline summarisation dataset Chen et al.
-(2023b).
-Thus we aim for the student to learn temporal reasoning and use this ability to generate timeline summaries.
+After we fine-tune a teacher model on the NarrativeReason dataset, we transfer the temporal reasoning knowledge to a student model, while, also fine-tuning the student on a news timeline summarisation dataset Chen et al.
+(2023).
+Thus we aim for the student to learn temporal reasoning and also use this ability when generating timeline summaries (step (b) in Fig.
+1).
 We fine-tune Phi-3-mini-4k-instruct as a student model.
 We use three knowledge distillation (KD) objectives to transfer knowledge from the teacher to the student: Neuron Selectivity Transfer (NST) Huang and Wang (2017), transfers heatmap like spatial activation patterns of teacher neurons to student neurons; Contrastive Representation Distillation (CRD) Tian et al.
-(2019), maximises the mutual information between the teacher and student representations with contrastive learning; Probabilistic Knowledge Transfer (PRT) Passalis and Tefas (2018), matches the probability distribution of the data in the feature space.
+(2019), maximises the mutual information between the teacher and student representations with contrastive learning; Probabilistic Knowledge Transfer (PRT) Passalis and Tefas (2018), matches the probability distribution of the data in the feature space of teacher and student models.
 PRT:
-The core idea is to match the probability distribution of the data in the feature space between teacher and student models.
-However, learning a significantly smaller model that accurately recreates the whole geometry of a complex teacher model is often impossible.
+Learning a significantly smaller model that accurately recreates the whole geometry of a complex teacher model is often impossible.
 Passalis and Tefas (2018) uses the conditional probability distribution to describe the samples.
-Here, $\mathbf{Y_{t}}={\{\mathbf{y_{t}}}_{1},{\mathbf{y_{t}}}_{2},...,{\mathbf{y_{t}}}_{l}\}\in\mathbb{R}^{vocab_{t}}$ denote the output logits of the teacher model, and $\mathbf{Y_{s}}={\{\mathbf{y_{s}}}_{1},{\mathbf{y_{s}}}_{2},...,{\mathbf{y_{s}}}_{l}\}\in\mathbb{R}^{vocab_{s}}$ denote the output logits of the student model, where ${\mathbf{y_{t}}}$/${\mathbf{y_{s}}}$ is vector and $l$ is the length of sentences, ${vocab_{t}}$ and ${vocab_{t}}$ are the vocabulary sizes of teacher and student models respectively.
+Here, $\mathbf{Y_{t}}={\{\mathbf{y_{t}}}_{1},{\mathbf{y_{t}}}_{2},...,{\mathbf{y_{t}}}_{l}\}\in\mathbb{R}^{vocab_{t}}$ denote the output logits of the teacher model, and $\mathbf{Y_{s}}={\{\mathbf{y_{s}}}_{1},{\mathbf{y_{s}}}_{2},...,{\mathbf{y_{s}}}_{l}\}\in\mathbb{R}^{vocab_{s}}$ denote the output logits of the student model, where ${\mathbf{y_{t}}}$ and ${\mathbf{y_{s}}}$ are vectors and $l$ is sentence length, ${vocab_{t}}$ and ${vocab_{s}}$ are the vocabulary sizes of teacher and student models respectively.
 We can define the conditional probability distribution for the teacher model as Eq.
 1, and student model as Eq.
-2.
+2, where $K$ is a symmetric kernel with scale parameter $\sigma$.
 ${p_{i|j}}=\frac{K({\mathbf{y_{t}}_{i}},\mathbf{y_{t}}_{j};2\sigma_{t}^{2})}{\sum_{k=1,k\neq{j}}^{l}{K(\mathbf{y_{t}}_{k},\mathbf{y_{t}}_{j};2\sigma_{t}^{2})}}$ (1)
-${q_{i|j}}=\frac{K(\mathbf{x_{s}}_{i},\mathbf{x_{s}}_{j};2\sigma_{s}^{2})}{\sum_{k=1,k\neq{j}}^{l}{K(\mathbf{x_{s}}_{k},\mathbf{x_{s}}_{j};2\sigma_{s}^{2})}}$ (2)
-In these two equations (Eq. 1 and Eq. 2), we use the cosine similarity as a kernel metric allows for more robust affinity estimations.
-${K_{consine}({\mathbf{y_{t}}_{i}},{\mathbf{y_{t}}_{j}})}=\frac{1}{2}(\frac{{\mathbf{y_{t}}_{i}^{\mathrm{T}}}\mathbf{y_{t}}_{j}}{\|{\mathbf{y_{t}}_{i}}\|{{}_{2}}\|{\mathbf{y_{t}}_{j}}\|{{}_{2}}}+1)\in[0,1].$
-We use Kullback-Leibler (KL) divergence to calculate the distance between the conditional probability distributions of the teacher model and the student model, as follows:
+${q_{i|j}}=\frac{K(\mathbf{y_{s}}_{i},\mathbf{y_{s}}_{j};2\sigma_{s}^{2})}{\sum_{k=1,k\neq{j}}^{l}{K(\mathbf{y_{s}}_{k},\mathbf{y_{s}}_{j};2\sigma_{s}^{2})}}$ (2)
+In equations Eq. 1 and Eq. 2, we use the cosine similarity kernel (no $\sigma$), allowing for more robust affinity estimations:
+${K_{cosine}({\mathbf{y_{t}}_{i}},{\mathbf{y_{t}}_{j}})}=\frac{1}{2}(\frac{{\mathbf{y_{t}}_{i}^{\mathrm{T}}}\mathbf{y_{t}}_{j}}{\|{\mathbf{y_{t}}_{i}}\|{{}_{2}}\|{\mathbf{y_{t}}_{j}}\|{{}_{2}}}+1)\in[0,1].$
+We use Kullback-Leibler (KL) divergence to calculate the distance between the conditional probability distributions of the teacher and student models:
 ${L_{PKT}}={\sum_{i=1}^{l}\sum_{j=1,i\neq{j}}^{l}}{p_{i|j}}\log({\frac{p_{i|j}}{q_{i|j}}}).$
@@
 Specifically, we transfer neuron selectivity knowledge from $\{{\mathbf{t(x)}}_{*,i}\}_{i=1}^{N}$ to $\{{\mathbf{s(x)}}_{*,i}\}_{i=1}^{M}$, where $N$ and $M$ are the hidden state dimensions.
-Then we follow the idea in Huang and Wang (2017) using Maximum Mean Discrepancy (MMD) to calculate the distance between these two distributions.
-The basic idea of MMD is that if the moments of two random variables are the same for all orders, then the two distributions are identical $L_{MMD^{2}}(\mathbf{t},\mathbf{s})$.
+Then we follow Huang and Wang (2017) in using Maximum Mean Discrepancy (MMD) to calculate the distance between the activation patterns of student $\{{\mathbf{s(x)}}_{*,i}\}_{i=1}^{M}$ and teacher neurons $\{{\mathbf{t(x)}}_{*,i}\}_{i=1}^{N}$.
+Here, we use squared MMD to calculate the diatance between $\mathbf{t}$ and $\mathbf{s}$.
 ${L_{MMD^{2}}(\mathbf{t},\mathbf{s})}={\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{i^{\prime}=1}^{N}K\left[{\mathbf{t(x)}}_{*,i};{\mathbf{t(x)}}_{*,i^{\prime}}\right]}$
 $+{{\frac{1}{M^{2}}\sum_{j=1}^{M}\sum_{j^{\prime}=1}^{M}K\left[{\mathbf{s(x)}}_{*,j};{\mathbf{s(x)}}_{*,j^{\prime}}\right]}}$
-$-{{\frac{1}{MN}\sum_{i=1}^{N}\sum_{j=1}^{M}k\left[{\mathbf{s(x)}}_{*,i};{\mathbf{t(x)}}_{*,j}\right]}}.$
-In this formula we use a Gaussian Kernel ${k(x,y)=exp(-\frac{{||x-y||}_{2}^{2}}{2{\sigma}^{2}})}$ with $\sigma=1$.
-We transfer the teacher activation patterns to the student by minimizing $L_{MMD^{2}}$.
+$-{{\frac{1}{MN}\sum_{i=1}^{N}\sum_{j=1}^{M}K\left[{\mathbf{s(x)}}_{*,i};{\mathbf{t(x)}}_{*,j}\right]}},$
+where ${K(x,y)=exp(-\frac{{||x-y||}_{2}^{2}}{2{\sigma}^{2}})}$ with $\sigma=1$ is Gaussian Kernel. We transfer the teacher activation patterns to the student by minimizing $L_{MMD^{2}}$.
 CRD
-maximises the lower-bound to the mutual information between the teacher and student representations.
-Here, we follow Tang et al.
-(2021) to sample one positive pair from the joint distribution $p(\mathbf{S},\mathbf{T})=q(\mathbf{S},\mathbf{T}|positive)$ for $N$ negative pairs sampled from the product of marginal $p(\mathbf{S})p(\mathbf{T})=q(\mathbf{S},\mathbf{T}|negtive)$, where $N$ is the batch size.
+maximizes the lower-bound to the mutual information between the teacher and student representations.
+In other words, we would like to push the representations $\mathbf{s(x}_{i})$ and $\mathbf{t(x}_{i})$ closer together, while pushing apart $\mathbf{s(x}_{i})$ and $\mathbf{t(x}_{j})$.
+Here, we follow the sampling process of Tang et al.
+(2021), providing 1 positive pair for every $N$ (batch size) negative pairs.
+The positive pair is sampled from the joint distribution $p(\mathbf{S},\mathbf{T})=q(\mathbf{S},\mathbf{T}|positive)$, and $N$ negative pairs are drawn from the product of marginals $p(\mathbf{S})p(\mathbf{T})=q(\mathbf{S},\mathbf{T}|negative)$, where $q$ is a distribution denoting whether the ($\mathbf{S}$,$\mathbf{T}$) pair is drawn from the positive or negative pairs.
 We can maximize the lower bound of mutual information by minimizing the following loss function:
@@
 $-N.\mathbb{E}_{q(\mathbf{s},\mathbf{t}|negtive)}\left[\log(1-h(\mathbf{s},\mathbf{t}))\right]$
-In Eq 3, function $h$ should satisfy $h:\{\mathbf{s},\mathbf{t}\}\rightarrow[0,1]$,
+In Eq 3, $h$ should satisfy $h:\{\mathbf{s},\mathbf{t}\}\rightarrow[0,1]$,
 ${h(\mathbf{s},\mathbf{t})}=\frac{exp({\mathbf{s}^{\mathrm{T}}}\mathbf{t})}{exp({\mathbf{s}^{\mathrm{T}}}\mathbf{t})+\frac{N}{M}},$
-where M is the cardinality of the dataset, and we need to normalize $\mathbf{s}$ and $\mathbf{t}$ by $L$-2 norm before the inner product.
-The knowledge distillation method transfers temporal reasoning knowledge from the teacher model to the student model.
-At the same time, we aim for this knowledge to benefit the timeline summarization task.
-Therefore, we fine-tune the student model on the timeline summarization dataset §4.1, enabling it to simultaneously learn from the teacher model and use the language modeling loss (for next token prediction) $L_{language}$ to integrate temporal reasoning knowledge with timeline summarization information.
+where M is the cardinality of the dataset, and we need to normalize $\mathbf{s}$ and $\mathbf{t}$ by $L$-2 norm before taking the inner product.
+The knowledge distillation process transfers temporal reasoning knowledge from the teacher to the student model.
+At the same time, we want this knowledge to benefit the timeline summarisation task.
+Thus, we fine-tune the student model on the timeline summarisation dataset §4.1, enabling it to both learn from the teacher model and use the language modeling loss (for next token prediction) $L_{language}$ to integrate temporal reasoning knowledge with timeline summarisation information.
 3.4 Mental Health Timeline Summary
@@
 We fine-tune the teacher model on the ‘NarrativeReason’ dataset.
-When distilling the temporal reason knowledge to the student model, we use a news timeline summarisation datasetChen et al.
-(2023a).
-Finally, we apply the model to a different domain, specifically generating mental health-related timeline summaries from social media.
+When distilling the temporal reasoning knowledge to the student model, we also fine-tune the latter on a news timeline summarisation dataset Chen et al.
+(2023).
+Finally, we apply the fine-tuned student model to a different domain, specifically that of generating mental health-related timeline summaries from social media.
 NarrativeReason We extracted 668 events from 30 articles, containing a total of 19,614 temporal relations between events Rogers et al.
@@
 We use these question/answer pairs to fine-tune the teacher model to enhance its temporal reasoning capability.
-Timeline summarisation Dataset This timeline summarisation dataset is sourced from Chen et al.
-(2023a).
-They collected timeline summaries from Wikipedia websites, with a total of 5,000 timelines and summaries.
+Timeline summarisation Dataset The timeline summarisation dataset used for training comes from Chen et al.
+(2023).
+It consists of timeline summaries from Wikipedia websites, with a total of 5,000 timelines and summaries.
+Due to some inconsistencies between events across timelines and summaries, this dataset was only used for training.
 TalkLife When generating the summary, we use the dataset collected by Tsakalidis et al.
-(2022) comprising 500 anonymised user timelines from Talklife.
+(2022) comprising 500 anonymised user timelines from Talklife  [^1] https://www.talklife.com/.
 Song et al.
-(2024) sample 30 timelines from it and annotate them with corresponding mental health-related summaries.
-These human-written summaries include diagnosis, intra- and interpersonal patterns and mental state changes over time, and they also highlight information related to individuals’ mental states, which can be used for automated evaluation.
+(2024) had sampled 30 timelines from it and augmented them with corresponding mental health-related summaries and associated evidence.
+The summaries cover aspects such as diagnosis, intra- and interpersonal patterns and mental state changes over time.
+They also highlighted associated evidence in the timelines, which they utilised for automated summary evaluation.
 4.2 Models & Baselines
-We compare our method against existing LLMs for mental health related summarisation. We introduce the implementation details in the appendix A.1
-L-phi This model uses LLaMA as the teacher model and a smaller model, Phi, as the student.
+For mental health related summarisation, we compare our method against existing LLMs . Implementation details are in appendix A.1
+L-Phi: This model derives from a LLaMA-3 teacher model teaching a smaller student model, Phi.
 We apply different knowledge distillation (KD) methods to transfer temporal reasoning knowledge to the student model.
+L-Phi is also fine-tuned for timeline summarisation.
 Subsequently, we directly prompt this model to generate mental health-related timeline summaries.
-P-phi In this model, we use Phi as the teacher model and another Phi of the same size as the student model.
-$\text{Phi}_{joint}$ To compare with the phi-phi, we use joint learning to fine-tune Phi on both the NarrativeReason and Timeline summarisation datasets.
-This allows us to observe whether knowledge distillation (KD) outperforms directly fine-tuning on the two datasets.
-In addition, we fine-tune Phi on the NarrativeReason and timeline summarisation datasets and obtain models $\text{Phi}_{temp}$ and $\text{Phi}_{timeline}$ separately.
-We use these two models to generate timeline summaries for comparison with the KD derived model.
-$\text{Phi}_{ICL}$ We use in context learning (ICL) to guide Phi to generate summaries.
+P-Phi: The same as L-Phi but we use another Phi as the teacher model instead of LLaMA-3.
+$\text{Phi}_{joint}$: To investigate the effect of knowledge distillation (KD), we compare P-Phi, against a joint learning setting where we fine-tune Phi on both the NarrativeReason and Timeline summarisation datasets.
+In this mixed dataset, we prepend a different prompt to clarify the task, ensuring that Phi can learn to apply its knowledge accordingly to each type of task.
+For details see appendix  A.3.
+$\text{Phi}_{temp}$ and $\text{Phi}_{tl}$: We fine-tune Phi on the NarrativeReason and timeline summarisation datasets separately and obtain models $\text{Phi}_{temp}$ and $\text{Phi}_{tl}$.
+We use these two models to generate timeline summaries for comparison, and examine whether fine-tuning on a single dataset can help the timeline summarisation task.
+$\text{Phi}_{ICL}$: We use in context learning (ICL) to guide Phi to generate summaries.
 We provide the model with a pair consisting of a timeline and its corresponding summary as an example, and then let it generate summaries for other timelines.
-LLaMA We prompt LLaMA to generate mental health related summaries on timelines directly as a baseline.
+LLaMA: We prompt LLaMA-3 (without any fine-tuning) to generate mental health related timeline summaries in a zero shot fashion.
+TH-VAE: We use TH-VAE Song et al. (2024) to generate mental health summaries for comparison, while we use LLaMA-3 to generate high-level summaries.
 4.3 Evaluation
-We use the timeline summaries from Song et al.
-(2024)4.1) forautomated evaluation.
-We employFactual Consistency (FC), to measure whether timeline summaries are consistent with the original timelines, and Evidence Appropriateness (EA), to measure the consistency of human written summaries with their corresponding timeline summaries Song et al.
+We work with the timeline summaries from Song et al.
+(2024)4.1).
+Like Song et al.
+(2024), we employ Factual Consistency (FC), to measure how consistent timeline summaries are with the original timelines, and Evidence Appropriateness (EA), to measure the consistency between human written summaries and corresponding timeline summaries Song et al.
 (2024).
-For human evaluation, we worked with two clinical psychology graduate students fluent in English to evaluate 30 summaries generated from 30 timelines (TalkLfe).
+Here, we use the annotated timeline evidence from Song et al.
+(2024) to directly generate high-level summaries.
+By contrast Song et al.
+(2024), generated clinically meaningful summaries from the timeline by first using TH-VAE to generate evidence related to mental health before generating high-level summaries.
+Since the goal of our paper is to validate the role of temporal reasoning in timeline summarisation, we directly use the annotated evidence to generate the high-level summary as the point is to make better use of the evidence rather than extract it from scratch.
+[^2] We obtained the annotated dataset from the authors Song et al.
+(2024).
+For human evaluation, we worked with two clinical psychology graduate students fluent in English to evaluate 30 summaries generated from 30 timelines (TalkLife).
 We follow the metrics used in Song et al.
 (2024) to evaluate the summaries from the perspectives of Factual Consistency and Usefulness (general/diagnosis/inter-&Interpersonal/MOC).
-A factually consistent summary should accurately represent the content of the timeline and excludes any information not present in the timeline.
-We also evaluate the usefulness of a mental health summary on the basis of four aspects: General (contains the most clinically important information from the timeline and aids the clinician in understanding a patient’s condition); Diagnosis (provides useful information about an individual’s diagnosis); inter-&Interpersonal (provides helpful information about an individuals’ main needs and patterns of self and other relationships); MOC (provides useful information about an individual’s changes over time in terms of emotion/cognition and behaviour).
+A factually consistent summary should accurately represent the content of the timeline.
+We also evaluate the quality of a mental health summary on the basis of: General usefulness (contains the most clinically important information from the timeline in understanding a patient’s condition); Diagnosis (provides useful information about an individual’s mental state); inter-&Interpersonal (provides helpful information about an individuals’ needs and relationship patterns); MOC (provides useful information about an individual’s changes over time ).
 5 Results and Discussion
-5.1 Automatic evaluation
-We conducted experiments with different combinations of KD methods.
-Table 1 shows experiment results.
-Since we did not change the output dimensions when fine-tuning LLaMA, the CRD method was not used during the KD process.
-Among the individual methods, PKT performed the best; however, combining PKT with NST achieved the best overall results.
+Automatic evaluation: We conducted experiments with different combinations of KD methods and fine-tuning strategies (Table 1).
+Since we didn’t change output dimensions when fine-tuning LLaMA, CRD was not used during the KD process.
+Among individual methods, PKT performed the best, but combining PKT with NST achieved the best overall results.
+P-Phi and L-Phi are the best performing models with identical EA (correspondence to human summaries).
+L-Phi seems to have lower FC, which would suggest it is less faithful to the original timeline.
+However, human evaluation in Table 2 indicates that clinical psychologists have a clear preference for summaries generated by L-Phi, on all aspects.
+This suggests distilling information from a larger LLM is indeed beneficial.
 […]
-Table 1: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) on student model.
-Higher is better, best in bold
-Table 2 shows the results of applying alternative fine-tuning strategies on LLMs.
-The results for $\text{Phi}_{temp}$ and $\text{Phi}_{timeline}$ on FC indicate that fine-tuning on a single dataset does not improve model performance; instead, it exacerbates hallucination issues.
-Notably, $\text{Phi}_{temp}$ performed the worst in both FC and EA metrics, suggesting that the incorporation of temporal reasoning information appears to interfere with the LLM’s ability to effectively handle the timeline summarization task.
-Additionally, in strategy $\text{Phi}_{jont}$, combining the two types of data directly during training failed to integrate them effectively.
-As a result, the performance of $\text{Phi}_{jont}$ was even worse than using in-context learning to guide the LLM ($\text{Phi}_{ICL}$).
-Metric $\text{Phi}_{ICL}$ $\text{Phi}_{temp}$ $\text{Phi}_{timeline}$ $\text{Phi}_{joint}$ FC .412 .141 .184 .238 EA .965 .895 .966 .941
-Table 2: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) across the compared models. Higher is better, best in bold
-5.2 Human evaluation
-Based on the results of the automatic evaluation, we selected the best-performing LLaMA-Phi and Phi-Phi models.
-Additionally, we included the non-fine-tuned versions of LLaMA and Phi.
+Table 1: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) for timeline summarisation of the different models.
+Higher is better.
+Best results are in bold and second-best results in underline.
+Regarding fine tuning strategies for timeline summarisation, the results for $\text{Phi}_{temp}$ and $\text{Phi}_{tl}$ on FC indicate that fine-tuning on a single dataset does not improve model performance; instead, it exacerbates hallucination issues.
+Notably, $\text{Phi}_{temp}$ performed the worst on both FC and EA metrics, suggesting that incorporating temporal reasoning information interferes with the LLM’s ability to effectively handle the timeline summarisation task.
+Additionally, in $\text{Phi}_{joint}$, combining the two types of data directly during training failed to integrate them effectively.
+As a result, the performance of $\text{Phi}_{joint}$ was worse than just using in-context learning to guide the LLM ($\text{Phi}_{ICL}$).
+Human evaluation: Based on the results of the automatic evaluation, we selected the best-performing L-Phi and P-Phi models.
+Additionally, we included the non-fine-tuned zero-shot versions of LLaMA and Phi.
 This can help us understand in which specific aspects the model has improved with the inclusion of temporal reasoning information.
-The fine-tuned model shows the greatest improvement in terms of factual consistency and usefulness (general).
+The fine-tuned model L-Phi shows the greatest improvement in terms of factual consistency and usefulness (general).
 This aligns with our findings when analyzing the summaries, where the fine-tuned model significantly reduces hallucination, as shown in Table 2.
 In addition, we found that the fine-tuned model did not show significant improvement in terms of Moments of Change (MOC).
-This could be because, in the timelines, users do not experience many emotional switches, resulting in the model-generated summaries being relatively similar.
-Aspect phi phi-phi LLaMA LLaMA-phi Factual Consistency 2.90 3.32 3.58 3.83 Usefulness (General) 2.60 3.13 3.17 3.48    (Diagnosis) 2.90 3.37 3.45 3.62   —-(Inter-& Intrapersonal) 2.95 3.00 3.40 3.51   —-(MoC) 2.97 2.97 3.42 3.47
+In Appendix  A.4, we give examples of summaries generated by different models, these examples clearly demonstrate that the fine-tuned model can effectively reduce hallucinations.
+Aspect Phi P-Phi LLaMA L-Phi Factual Consistency 2.90 3.32 3.58 3.83 Usefulness (General) 2.60 3.13 3.17 3.48    (Diagnosis) 2.90 3.37 3.45 3.62   —-(Inter-& Intrapersonal) 2.95 3.00 3.40 3.51   —-(MoC) 2.97 2.97 3.42 3.47
 Table 2: Human evaluation results based on 5-point Likert scales (1 is worst, 5 is best). Best in bold.
 5.1 Why knowledge distillation works
-In this section, we analyze why the $\text{Phi}_{timeline}$ model performed better from a representation learning perspective.
-For this purpose, we perform two experiments, a task understanding probing experiment and a Joint Task Representation Learning (JTRL) experiment.
-We analyze the internal representations of $\text{Phi}_{timeline}$ against $\text{Phi}_{joint}$.
-We construct our probing dataset by using $\text{Phi}_{joint}$’s test dataset.
-We pose the two tasks as a binary classification problem and then extract the activations for the last layer.
-Then we use UMAP (McInnes et al., 2018) to project these activations to lower dimensions (Sainburg et al., 2021; Tseriotou et al., 2023).
-From Figure 3, we can see that the activations of $\text{Phi}_{joint}$ are well separated for each task, whereas the activations for $\text{Phi}_{timeline}$ overlap.
-Given the performance of $\text{Phi}_{timeline}$, this leads us to hypothesize that the model learned better representations for the task due to more task specific polysemantic (Olah et al., 2020) neurons.
-To validate our hypothesis, we ran another set of experiments to analyze the internal representation difference between the two models, which we termed as JTRL.
-Our main idea here was to validate that knowledge distillation resulted in better representations due to more polysemantic neurons, and these representations varied highly compared to $\text{Phi}_{joint}$’s representation.
+[…]
+To validate our hypothesis, we ran another set of experiments (JTRL) to analyze the internal representation difference between the two models.
+Our hypothesis here is that knowledge distillation results in better representations due to more polysemantic neurons, whose representations vary significantly to those of $\text{Phi}_{joint}$.
 To measure JTRL, we use the Centered Kernel Alignment (CKA) (Kornblith et al., 2019) similarity score.
@@
 Then, we calculate the CKA similarity score between the mean sentence embeddings and each layer representation of the model.
-We did this both for the individual models and between the layers of the models.
-From Figure 4, we can see that $\text{Phi}_{timeline}$ shows a gradual increase in CKA across layers, peaking in the mid-to-late layers.
-This indicates that as the layers progress, $\text{Phi}_{timeline}$ preserves and refines the information from the initial embeddings in a task-relevant way.
-The is likely due to the fact that the distillation process encourages this alignment by transferring task-relevant knowledge from the teacher.
-On the other hand, $\text{Phi}_{joint}$ shows an initial increase but saturates and flattens early.
-This indicates that it fails to refine representations effectively in deeper layers, which could be due to conflicting objectives between the tasks.
-The lower CKA in later layers suggests that the model moves away from the initial embeddings in a manner that is less effective for task-specific learning.
-Lastly, the CKA values between the models show that they learn vastly different representations.
-In short, $\text{Phi}_{timeline}$’s ability to align with the initial embeddings correlates with its better task performance, as the CKA value reflects how well the model retains and transforms meaningful input information throughout its layers.
+We did this both for the individual models and between models’ layers.
+From Figure 4, we can see that $\text{Phi}_{tl}$ shows a gradual increase in CKA across layers, peaking in the mid-to-late layers.
+This indicates that as the layers progress, $\text{Phi}_{tl}$ preserves and refines the information from the initial embeddings in a task-relevant way.
+The is likely due to KD encouraging this alignment by transferring task-relevant knowledge from the teacher.
+While $\text{Phi}_{joint}$ shows an initial increase in CKA, it saturates and flattens early.
+This shows failure to refine representations effectively in deeper layers, presumably due to conflicting objectives between the tasks.
+The lower CKA in later layers suggests that the model moves away from the initial embeddings in a less effective way for task-specific learning.
+Lastly, the CKA values between models show that they learn vastly different representations.
+In short, $\text{Phi}_{tl}$’s ability to align with the initial embeddings correlates with its better task performance, as the CKA value reflects how well the model retains and transforms meaningful input information throughout its layers.
 (a) UMAP projection $\text{Phi}_{joint}$
-(b) UMAP projection $\text{Phi}_{timeline}$
+(b) UMAP projection $\text{Phi}_{tl}$
 […]
-Figure 4: CKA similarity score of both within and between $\text{Phi}_{timeline}$ and $\text{Phi}_{joint}$ model representations.
+Figure 4: CKA similarity score of both within and between $\text{Phi}_{tl}$ and $\text{Phi}_{joint}$ model representations.
 6 Conclusions
-We created a dataset named NarrativeReason, which is utilised to enhance the temporal reasoning abilities of LLMs.
-This dataset enables models to uncover temporal relationships between events within timelines.
-We fine-tune a larger LLM with NarrativeReason, then distill its enhanced temporal reasoning capability to a smaller LLM while leveraging the distilled knowledge to enhance performance in timeline summarisation tasks.
+We have created a dataset (NarrativeReason), containing relation triples to represent event sequences within narratives.
+We fine-tune a large LLM on NarrativeReason, then use knowledge distillation (KD) to transfer its enhanced temporal reasoning capability to a smaller LLM while leveraging the distilled knowledge to enhance performance in timeline summarisation tasks.
 We apply the model to a different domain from the one it was trained on, namely generating mental health related timeline summaries.
-Our results demonstrate that our approach using knowledge distillation produces more accurate summaries while significantly reducing hallucinations.
-We further analyzed the internal representations of the model and found that knowledge distillation leads to better feature representations within the model that are more aligned with the task of timeline summarisation.
+Our results demonstrate that our KD approach produces more accurate summaries while significantly reducing hallucinations.
+Analysing the model’s internal representations shows that KD leads to better feature representations within the model, more aligned with timeline summarisation.
 Limitations
@@
 A.2 Experiment on teacher model
-After fine-tuning the teacher model, we used the TEMPREASON dataset Tan et al.
-(2023b) to evaluate its performance pre- and post-fine-tuning.
+After fine-tuning the teacher model, we used the TEMPREASON dataset tan2023towards to evaluate its performance pre- and post-fine-tuning.
 Specifically, we focused on the L3 part of the dataset, which deals with event-event relations, to determine whether the model could accurately infer the temporal sequence between two events.
 We use F1 as the evaluation metrics, the experimental results show that the fine-tuned model achieved a 0.07 improvement in this metric compared to the pre-fine-tuning model, highlighting its enhanced ability to infer event-event temporal relations.
+A.3 Joint learning
+When fine-tuning Phi using joint learning, we mix the timeline summarization and temporal reasoning datasets and shuffle their order.
+We then fine-tune Phi using this mixed dataset.
+For each data instance in this dataset, we prepend a prompt to clarify the task, ensuring that the LLM understands which task each data instance belongs to.
+For example, for timeline summarization data, we add the prompt "This is a timeline summarization task, your task is to summarize the provided timeline + timeline summarisation text".
+Similarly, for temporal reasoning tasks, we would prepend a different prompt to clarify the task at hand, ensuring that Phi can learn to apply its knowledge accordingly to each type of task.
+A.4 Examples of summaries generated in different models
+Gold (Diagnosis)This teenager girl had attempted suicide and is at high risk of trying again.
+She is depressed and lonely.
+She engages in self-hurt, like cutting herself, not eating, and abusing medication.
+She is being bullied at school, physically and sexually harassed.
+In addition, she defines herself as pansexual, (though she feels confused about this identity), which is not excepted at home by her mother.
+She also fears not being accepted by her friends.
+Furthermore, she has difficulty sleeping and suffers from nightmares.
+The home environment is stressful.
+Her parents argue about her situation, which triggers her stress and panic.
+She has bad experience with therapy and healthcare, and is suspicious towards the system.
+Her sense of agency is low.
+(IIP)She is longing for love and companionship.
+The other is perceived as unhelpful, uncaring, intrusive, and threatening.
+Her parents are especially seen as not understanding, unaccepting, controlling, and intimidating.
+The response of self is self-pity, self-blame, and hopelessness.
+(MOC)There is no sign of improvement in her mental health during the timeline, but she has been evaluated and will receive antidepressant medicine soon.
+L-Phi (Diagnosis)The individual seems to be grappling with significant emotional turmoil, likely stemming from depression or anxiety disorders.
+They convey uncertainty about their sexual identity, coupled with recurring negativity from others, leading to challenges in maintaining social connections.
+Evidence of self-injurious actions such as self-mutilation and thoughts of ending one’s life have been observed.
+This person has faced relentless torment within educational settings, intensifying their alienation and despondency.
+Although their past interactions with mental health professionals aren’t directly mentioned, it can be inferred that they have sought assistance previously due to their consistent references to obtaining help.
+Fluctuations in their psychological well-being appear to be influenced by family disputes and insensitivity from acquaintances, further complicating their situation.
+(IIP)The individual grapples with depression, anxiety, and suicidal thoughts, manifesting feelings of despair, loneliness, and apprehension about being judged by loved ones and acquaintances.
+These experiences highlight their urgent need for empathy and comprehension from those around them, yet they find themselves subjected to criticism and misunderstanding owing to their sexual orientation and mental wellbes.
+Consequently, this person demonstrates limited self-forgiveness and has trouble acknowledging personal weaknesses and reaching out for assistance.
+(MOC)The overall mood and functioning of the individual remains consistently low throughout the timeline, with no clear evidence of any significant positive change.
+Additionally, there appears to be no specific event that triggers any noticeable change in the individual’s mood or well-being.
+P-Phi (Diagnosis) Based on the provided social media posts, the individual appears to be struggling with several mental health challenges including depression, anxiety, self-harm tendencies, and possible substance misuse.
+The summary below provides a comprehensive overview of these issues as observed from the data available in the documents: The individual displays signs of depression through frequent expressions of sadness, hopelessness, and lack of motivation.
+Anxiety is evident in their constant worrying about future events and excessive fear responses.
+Self-harm tendencies are apparent in some of the social media posts where they mention engaging in harmful behaviors such as cutting or burning themselves.
+Lastly, there are indications of possible substance misuse, as they frequently post about using drugs or alcohol to cope with their emotional struggles.
+Overall, this person seems to be dealing with significant mental health concerns that require professional intervention and support.
+(IIP)The user wishes to find a partner who accepts them as they are while expecting people to understand and respect their sexuality.
+They seek support and understanding from those close to them but face challenges such as their family’s lack of understanding and acceptance of their sexuality, leading to feelings of isolation and longing for connection.
+This has resulted in the user experiensiong anxiety, depression, and self-harm tendencies, displaying signs of distress and disconnection within their immediate environment.
+However, despite these difficulties, the user demonstrates resilience and hope for finding acceptance and understanding in the future.", (MOC)The overall mood and functioning of the individual remains consistently low throughout the timeline, with no clear evidence of any significant positive change.
+Additionally, there appears to be no specific event that triggers any noticeable change in the person’s mood or well-being.
+LLaMA (Diagnosis) This individual presents with severe depression, characterized by persistent suicidal ideation, feelings of worthlessness, and hopelessness.
+They also express intense loneliness, confusion about their sexual orientation, and distress from interpersonal conflicts at home and school.
+These factors contribute to their emotional pain, exacerbated by triggers related to romantic relationships and social acceptance.
+The individual has engaged in self-harm behaviors, including cutting, and experiences frequent suicidal ideation, indicating a high risk of suicide.
+They have shown some awareness of the need for professional help, having sought therapy and started anti-depressant medication, although they currently lack regular therapeutic support due to their therapist’s absence.
+Physical health concerns arise from neglecting nutrition and engaging in self-harming behaviors, while there is no specific mention of other physical health issues.
+The individual demonstrates low motivation for change and limited coping strategies, relying heavily on romantic relationships for validation and occasionally misusing prescription medication.
+Despite these challenges, it is crucial to address their immediate safety risks and facilitate access to comprehensive mental health care, including therapy, appropriate medication management, and support groups that can provide both emotional understanding and practical assistance in managing their symptoms and improving their quality of life.
+(IIP)<<ANALYSIS>> The individual displays a deep-seated wish for a romantic partnership, frequently expressing a desire for a g̈irlfriend or boyfriend.T̈hey grapple with their sexual orientation, identifying as potentially pansexual but expressing uncertainty about their own understanding.
+Their pressing need for companionship and support is apparent, highlighted by their loneliness and fear of revealing their true self to friends.
+Moreover, they face considerable mental health challenges, including suicidal ideation, self-harming behaviors, and difficulty managing their prescribed medications.
+Interactions with peers are perceived negatively, with the individual reporting bullying and feeling victimized rather than supported.
+There is no evidence of recognizing complexity in others’ behavior; instead, social exchanges are colored by negativity and conflict.
+Their response to others is marked by fear, vulnerability, and a pervasive sense of being misunderstood.
+The fear of coming out to friends due to potential rejection or negative reactions underscores a profound lack of confidence in interpersonal relations and a heightened sensitivity to judgment from others.
+On a deeply personal level, the individual experiences overwhelming despair and self-blame.
+They consistently berate themselves (Ï’m such a screw-up)̈, burdened by guilt (Ït’s all my fault)̈, reflecting a severe erosion of self-esteem and self-worth.
+This self-criticism is coupled with a high risk of self-harm and persistent thoughts of suicide, demonstrating a pronounced inability to practice self-compassion and accept their own strengths and weaknesses.
+A dysfunctional intrapersonal pattern.
+(MOC)The individual’s emotional state fluctuates significantly between periods of profound despair and intermittent moments of hopefulness.
+These shifts are closely tied to various stressors, including family conflicts, school-related issues, and personal identity struggles.
+Phi (Diagnosis)As an AI developed by Microsoft, I am programmed to follow strict ethical guidelines prioritizing user safety and privacy.
+Consequently, I can only analyze personal data in a secure and consensual environment specifically tailored for professional use.
+In a hypothetical scenario where such conditions are met, a trained professional would cautiously assess the situation, taking care to uphold confidentiality and abide by legal standards.
+This case involves an individual grappling with various mental health issues, including depression, anxiety, potential self-harm behaviors, and profound despair.
+Their struggle extends to uncertainties about sexual orientation, experiences of bullying, familial discord, and negative remarks stemming from disclosed private information, all contributing to significant distress over body image and self-worth.
+Alarmingly, they have openly expressed thoughts of suicide, underscoring the critical need for immediate intervention by certified mental health experts.
+It must be emphasized that this description serves as an example and does not substitute for actual therapeutic engagement with a licensed practitioner.
+(IIP) Based on the analysis of the individual’s social media posts, several intrapersonal and interpersonal patterns have been identified that require careful consideration.
+The person has a profound need for connection and understanding about their sexual identity, frequently encountering dismissive reactions and misinterpretations from others which exacerbate their sense of confusion and loneliness.
+Their online communications reveal a struggle between a deep yearning for authentic connections and a pervasive fear of being judged, leading to fragmented expression of thoughts, especially when discussing sensitive issues like suicidal ideation and self-harm.
+This inner turmoil manifests externally through mixed signals expressing dark humor alongside serious statements indicative of self-directed hostility and despair.
+Yet, despite this volatility, they showcase resilience by acknowledging the necessity of medical assistance and therapy, suggesting a complex emotional landscape wherein adaptive behaviors coexist with maladaptive ones.
+Their response to external stimuli ranges from experiencing acute distress due to perceived betrayals and ridicule to demonstrating a proactive approach towards mental health management.
+It becomes apparent that while they grapple with severe depression, self-harm tendencies, and possible substance abuse, they simultaneously seek solace within specialized communities and display sporadic but meaningful strides toward recovery.
+Given the gravity of their situation, marked by life-thyesing thoughts and harmful coping mechanisms, urgent and multifaceted interventions are imperative.
+Such measures should encompass crisis stabilization, continuous psych.
+(MOC)The individual’s mood, well-being, and functioning exhibit significant fluctuations over time, transitioning between periods of despair and suicidal ideation as well as moments when they seek connection and support.
+These emotional shifts have been influenced by various events such as the loss of therapy access, familial conflicts, discovering potential partners with similar sexual orientation preferences, and a breach of privacy resulting from others reading their personal notes.
+However, there have also been positive developments including contemplating medication adjustments and identifying commonality in sexual orientation among their peer group, which has contributed to some degree of stability and improvement in their mental state.
+Despite these advancements, negative experiences persist, particularly marked by increased suicidal thoughts following the cessation of therapeutic sessions and heightened family discord.
+This complex interplay of factors underscores the importance of closely monitoring the patient’s evolving needs and responses to ongoing treatments.
+Table 3: We have highlighted hallucinations and obvious errors in yellow.
+Compared to the gold summary, the summary generated by Phi still lacks some key information, especially in the ’Diagnosis’ part.

### v3

@@
 This paper explores whether enhancing temporal reasoning capabilities in Large Language Models (LLMs) can improve the quality of timeline summarisation, the task of summarising long texts containing sequences of events, such as social media threads.
-We first introduce NarrativeReason, a novel dataset focused on temporal relationships among sequential events within narratives, distinguishing it from existing temporal reasoning datasets that primarily address pair-wise event relationships.
+We first introduce NarrativeReason, a novel dataset focused on temporal relations among sequential events within narratives, distinguishing it from existing temporal reasoning datasets that primarily address pair-wise event relations.
 Our approach then combines temporal reasoning with timeline summarisation through a knowledge distillation framework, where we first fine-tune a teacher model on temporal reasoning tasks and then distill this knowledge into a student model while simultaneously training it for the task of timeline summarisation.
@@
 This raises the question of whether temporal reasoning can be leveraged to enhance the quality of complex timeline summaries.
-Temporal reasoning involves understanding and processing temporal information in text to deduce time-based relationships between events Wenzel and Jatowt (2023a).
+Temporal reasoning involves understanding and processing temporal information in text to deduce time-based relations between events Wenzel and Jatowt (2023a).
 Zhou et al.
@@
 (2023); Chu et al.
-(2024); Wang and Zhao (2024), NarrativeReason focuses on the temporal relationships among a series of events within a story rather than distinct event pairs.
+(2024); Wang and Zhao (2024), NarrativeReason focuses on the temporal relations among a series of events within a story rather than distinct event pairs.
 This can help LLMs process a series of events to generate a coherent and accurate timeline summary.
@@
 (2024).
-Not only does it generate more accurate summaries but it also reduces hallucinations in LLMs.
+Not only does it generate more accurate summaries, but it also reduces hallucinations in LLMs.
   • $\bullet$
@@
 (2024) use an aligned timeline to improve an LLM’s temporal reasoning by translating the context into a temporal graph,identifying valid time expressions and generating related temporal knowledge.
-The temporal relationship between events is inferred based on specific times (e.g., year of event).
+The temporal relations between events are inferred based on specific times (e.g., year of event).
 However in a narrative, events often occur without a clear indication of time.
@@
 (2011b) developed a natural language understanding (NLU) system with a temporal reasoning component to create comprehensive timelines, applied to medical records, presenting medical history in a more intuitive way.
-They found that temporal reasoning in NLU is tightly integrated into the NLP system’s deep semantic analysis and can help a LM analyze temporal relationships between different events, which is beneficial for event or news summarisation Vashishtha et al.
+They found that temporal reasoning in NLU is tightly integrated into the NLP system’s deep semantic analysis and can help a LM analyze temporal relations between different events, which is beneficial for event or news summarisation Vashishtha et al.
 (2020).
@@
 (2024), which in turn had re-annotated TimeBankDense Cassidy et al.
-(2014) with a timeline-based annotation framework, NarrativeTime framework.
+(2014) with a timeline-based annotation framework, NarrativeTime.
 Rogers et al.
-(2024) had annotated all possible temporal links (TLINKS) between all events occurring within a narrative, thus providing temporal relations for the entire sequence of events (timeline) rather than just between event pairs to get a NarrativeTime dataset.
+(2024) have annotated all possible temporal links (TLINKS) between all events occurring within a narrative, thus providing temporal relations for the entire sequence of events (timeline) rather than just between event pairs to get obtain the NarrativeTime dataset.
 Here, TLINKs usually denote the event order information (e.g., before, after, during).
@@
 (2003), linking the corresponding arguments.
-However, NarrativeTime dataset  Rogers et al.
-(2024) only denotes the type of the event, annotated at the level of the verb, e.g.
-for the sentence "the value of the Indonesian stock market has fallen by twelve percent" this is annotated in NarrativeTime datasets as "the value of the Indonesian stock market has <EVENT class="OCCURRENCE" eid="e7"> fallen </EVENT> by twelve percent".
+However, NarrativeTime  Rogers et al.
+(2024) only denotes the type of event, annotated at the level of the verb, e.g.
+for the sentence "the value of the Indonesian stock market has fallen by twelve percent" this is annotated in the NarrativeTime dataset as "the value of the Indonesian stock market has <EVENT class="OCCURRENCE" eid="e7"> fallen </EVENT> by twelve percent".
 However, this only constitutes a denser temporal annotation without providing event triples and is therefore unsuited for temporal reasoning training tasks.
@@
 In order to represent an event, we use verbs as triggers to extract relational triples e.g.
-<Indonesian stock market value, fallen, by twelve percent> to represent the event annotated in NarrativeTime dataset as fallen.
-Then, we use these triples to construct the temporal relations of a series event.
+<Indonesian stock market value, fallen, by twelve percent> now represents the event annotated in the NarrativeTime dataset as fallen.
+Then, we use these triples to construct the temporal relations of a series of events.
 (e.g.
@@
 (2022).
-The input and output of the model are the temporal questions and corresponding answers.
+The input and output of the model are the temporal questions and corresponding answers respectively.
 Our experiments show that indeed fine-tuning on NarrativeReason improves performance on established temporal reasoning tasks (Appendix A.2) .
 Figure 2: The temporal relations between events.
-The text in the left column comes from NarrativeTime dataset, and we filtered it to keep only verbs.
+The text in the left column comes from the NarrativeTime dataset, and we filtered it to keep only verbs.
 We represent events triggered by verbs using relational triples, as shown in the middle column.
@@
 $-{{\frac{1}{MN}\sum_{i=1}^{N}\sum_{j=1}^{M}K\left[{\mathbf{s(x)}}_{*,i};{\mathbf{t(x)}}_{*,j}\right]}},$
-where ${K(x,y)=exp(-\frac{{||x-y||}_{2}^{2}}{2{\sigma}^{2}})}$ with $\sigma=1$ is Gaussian Kernel. We transfer the teacher activation patterns to the student by minimizing $L_{MMD^{2}}$.
+where ${K(x,y)=exp(-\frac{{||x-y||}_{2}^{2}}{2{\sigma}^{2}})}$ with $\sigma=1$ is the Gaussian Kernel. We transfer the teacher activation patterns to the student by minimizing $L_{MMD^{2}}$.
 CRD
@@
 At the same time, we want this knowledge to benefit the timeline summarisation task.
-Thus, we fine-tune the student model on the timeline summarisation dataset §4.1, enabling it to both learn from the teacher model and use the language modeling loss (for next token prediction) $L_{language}$ to integrate temporal reasoning knowledge with timeline summarisation information.
+Thus, we fine-tune the student model on the timeline summarisation dataset §4.1, enabling it to both learn from the teacher model and use the language modeling loss $L_{language}$ (for next token prediction) to integrate temporal reasoning knowledge with timeline summarisation information.
 3.4 Mental Health Timeline Summary
 We apply the student model to other domains, specifically to generate mental health-related summaries for timelines §4.1 from social media.
-For the mental health summary, we use the format proposed by Song et al.
+For mental health summaries, we use the format proposed by Song et al.
 (2024), which includes three key clinical concepts (diagnosis, inter- and intra- personal relations, moments of change).
@@
 Finally, we apply the fine-tuned student model to a different domain, specifically that of generating mental health-related timeline summaries from social media.
-NarrativeReason We extracted 668 events from 30 articles, containing a total of 19,614 temporal relations between events Rogers et al.
+NarrativeReason We extracted 668 events from 30 articles, containing a total of 19,614 temporal relations between events in Rogers et al.
 (2024), leading to 19,614 question/answer pairs for event-event relations.
@@
 Subsequently, we directly prompt this model to generate mental health-related timeline summaries.
-P-Phi: The same as L-Phi but we use another Phi as the teacher model instead of LLaMA-3.
+P-Phi: The same configuration as L-Phi but we use another Phi as the teacher model instead of LLaMA-3.
 $\text{Phi}_{joint}$: To investigate the effect of knowledge distillation (KD), we compare P-Phi, against a joint learning setting where we fine-tune Phi on both the NarrativeReason and Timeline summarisation datasets.
@@
 We provide the model with a pair consisting of a timeline and its corresponding summary as an example, and then let it generate summaries for other timelines.
+$\text{KD}_{timeline}$: To assess whether temporal reasoning knowledge contributes to the observed positive improvements, we fine-tune the teacher with the timeline summarisation dataset, and transfer it to the student model.
+$\text{KD}_{origin}$: We also transfer knowledge from the teacher model without finetuning on any dataset.
+Instead, we examine whether the teacher’s inherent knowledge can help improve the timeline summarization task.
 LLaMA: We prompt LLaMA-3 (without any fine-tuning) to generate mental health related timeline summaries in a zero shot fashion.
-TH-VAE: We use TH-VAE Song et al. (2024) to generate mental health summaries for comparison, while we use LLaMA-3 to generate high-level summaries.
+TH-VAE: For comparison with the state-of-the-art on this dataset we use TH-VAE Song et al.
+(2024) to generate mental health timeline summaries which are then translated by LLaMA-3 to high-level summaries.
 4.3 Evaluation
@@
 (2024).
+FC and EA combined effectively capture whether the generated summary aligns with factual information.
+Thus increase in these scores means reduction in hallucinations and generation of more reliable summaries.
 Here, we use the annotated timeline evidence from Song et al.
@@
 (2024), generated clinically meaningful summaries from the timeline by first using TH-VAE to generate evidence related to mental health before generating high-level summaries.
-Since the goal of our paper is to validate the role of temporal reasoning in timeline summarisation, we directly use the annotated evidence to generate the high-level summary as the point is to make better use of the evidence rather than extract it from scratch.
-[^2] We obtained the annotated dataset from the authors Song et al.
-(2024).
+Since the goal of our paper is to validate the role of temporal reasoning in timeline summarisation, we directly use the annotated evidence in the dataset to generate the high-level summary as the point is to make better use of the evidence rather than extract it from scratch.
 For human evaluation, we worked with two clinical psychology graduate students fluent in English to evaluate 30 summaries generated from 30 timelines (TalkLife).
@@
 5 Results and Discussion
-Automatic evaluation: We conducted experiments with different combinations of KD methods and fine-tuning strategies (Table 1).
+Automatic evaluation: We conducted experiments with different combinations of KD methods (Table 1).
 Since we didn’t change output dimensions when fine-tuning LLaMA, CRD was not used during the KD process.
 Among individual methods, PKT performed the best, but combining PKT with NST achieved the best overall results.
+We know that NST matches the distributions of neuron selectivity patterns between teacher and student networks; PRT also matches the probability distribution of the data in the feature space of teacher and student models.
+However, CRD focuses on distinguishing between different instances rather than learning structural consistency.
+In learning temporal reasoning, contrastive learning (CRD) may not be as effective as NST and PRT, since the task requires maintaining temporal coherence rather than separating instances.
 P-Phi and L-Phi are the best performing models with identical EA (correspondence to human summaries).
@@
 […]
-Table 1: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) for timeline summarisation of the different models.
+Table 1: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) for timeline summarisation of the different KD strategies.
 Higher is better.
 Best results are in bold and second-best results in underline.
-Regarding fine tuning strategies for timeline summarisation, the results for $\text{Phi}_{temp}$ and $\text{Phi}_{tl}$ on FC indicate that fine-tuning on a single dataset does not improve model performance; instead, it exacerbates hallucination issues.
+[…]
+Table 2: Automatic evaluation for factual consistency (FC), evidence appropriateness (EA) for timeline summarisation of all the different models we consider.
+Higher is better.
+Best results are in bold and second-best results in underline.
+Regarding fine tuning strategies for timeline summarisation as shown in Table 2 , the results for $\text{Phi}_{temp}$ and $\text{Phi}_{tl}$ on FC indicate that fine-tuning on a single dataset does not improve model performance; instead, it exacerbates hallucination issues.
 Notably, $\text{Phi}_{temp}$ performed the worst on both FC and EA metrics, suggesting that incorporating temporal reasoning information interferes with the LLM’s ability to effectively handle the timeline summarisation task.
-Additionally, in $\text{Phi}_{joint}$, combining the two types of data directly during training failed to integrate them effectively.
+Additionally, in $\text{Phi}_{joint}$, combining the two types of data directly during training, failed to integrate them effectively.
 As a result, the performance of $\text{Phi}_{joint}$ was worse than just using in-context learning to guide the LLM ($\text{Phi}_{ICL}$).
+[…]
 Human evaluation: Based on the results of the automatic evaluation, we selected the best-performing L-Phi and P-Phi models.
@@
 In addition, we found that the fine-tuned model did not show significant improvement in terms of Moments of Change (MOC).
-In Appendix  A.4, we give examples of summaries generated by different models, these examples clearly demonstrate that the fine-tuned model can effectively reduce hallucinations.
+In Appendix  A.4, we give examples of summaries generated by different models.
+These examples clearly demonstrate that the fine-tuned model can effectively reduce hallucinations.
 Aspect Phi P-Phi LLaMA L-Phi Factual Consistency 2.90 3.32 3.58 3.83 Usefulness (General) 2.60 3.13 3.17 3.48    (Diagnosis) 2.90 3.37 3.45 3.62   —-(Inter-& Intrapersonal) 2.95 3.00 3.40 3.51   —-(MoC) 2.97 2.97 3.42 3.47
@@
 5.1 Why knowledge distillation works
-[…]
+Here we analyze from a representation learning perspective why the L-Phi model performs better.
+We run two experiments on: (1)task understanding probing experiment and (2) Joint Task Representation Learning (JTRL).
+We analyze the internal representations of L-Phi against $\text{Phi}_{joint}$.
+We construct our probing dataset from the test dataset of the latter.
+We pose the two tasks as binary classification and extract activations for the last layer.
+We use UMAP (McInnes et al., 2018) to project the activations to lower dimensions (Sainburg et al., 2021; Tseriotou et al., 2023).
+In Figure 3, we can see that the activations of $\text{Phi}_{joint}$ are well separated for each task, whereas those of L-Phi overlap.
+Given the performance of L-Phi, this would suggest that the model learned better representations for the tasks due to more task-specific polysemantic (Olah et al., 2020) neurons.
 To validate our hypothesis, we ran another set of experiments (JTRL) to analyze the internal representation difference between the two models.
@@
 We did this both for the individual models and between models’ layers.
-From Figure 4, we can see that $\text{Phi}_{tl}$ shows a gradual increase in CKA across layers, peaking in the mid-to-late layers.
-This indicates that as the layers progress, $\text{Phi}_{tl}$ preserves and refines the information from the initial embeddings in a task-relevant way.
+From Figure 4, we can see that L-Phi shows a gradual increase in CKA across layers, peaking in the mid-to-late layers.
+This indicates that as the layers progress, L-Phi preserves and refines the information from the initial embeddings in a task-relevant way.
 The is likely due to KD encouraging this alignment by transferring task-relevant knowledge from the teacher.
@@
 Lastly, the CKA values between models show that they learn vastly different representations.
-In short, $\text{Phi}_{tl}$’s ability to align with the initial embeddings correlates with its better task performance, as the CKA value reflects how well the model retains and transforms meaningful input information throughout its layers.
+In short, L-Phi’s ability to align with the initial embeddings correlates with its better task performance, as the CKA value reflects how well the model retains and transforms meaningful input information throughout its layers.
 (a) UMAP projection $\text{Phi}_{joint}$
-(b) UMAP projection $\text{Phi}_{tl}$
+(b) UMAP projection L-Phi
 […]
-Figure 4: CKA similarity score of both within and between $\text{Phi}_{tl}$ and $\text{Phi}_{joint}$ model representations.
+Figure 4: CKA similarity score of both within and between L-Phi and $\text{Phi}_{joint}$ model representations.
 6 Conclusions
@@
 Significant efforts are required to establish the scientific validity of the clinical benefits offered by these summaries before they can be integrated into routine clinical practice.
+Acknowledgments
+This work was supported by a UKRI/EPSRC Turing AI Fellowship to Maria Liakata (grant ref EP/V030302/1) and the Alan Turing Institute (grant ref EP/N510129/1).
+This work was also supported by the Engineering and Physical Sciences Research Council [grant number EP/Y009800/1], through funding from Responsible Ai UK (KP0016) as a Keystone project lead by Maria Liakata.
 Appendix A Appendix
@@
 A.2 Experiment on teacher model
-After fine-tuning the teacher model, we used the TEMPREASON dataset tan2023towards to evaluate its performance pre- and post-fine-tuning.
+After fine-tuning the teacher model, we used the TEMPREASON dataset Tan et al.
+(2023) to evaluate its performance pre- and post-fine-tuning.
 Specifically, we focused on the L3 part of the dataset, which deals with event-event relations, to determine whether the model could accurately infer the temporal sequence between two events.
@@
 A.3 Joint learning
-When fine-tuning Phi using joint learning, we mix the timeline summarization and temporal reasoning datasets and shuffle their order.
-We then fine-tune Phi using this mixed dataset.
-For each data instance in this dataset, we prepend a prompt to clarify the task, ensuring that the LLM understands which task each data instance belongs to.
-For example, for timeline summarization data, we add the prompt "This is a timeline summarization task, your task is to summarize the provided timeline + timeline summarisation text".
-Similarly, for temporal reasoning tasks, we would prepend a different prompt to clarify the task at hand, ensuring that Phi can learn to apply its knowledge accordingly to each type of task.
+[…]
 A.4 Examples of summaries generated in different models